### v1

# Generalized Frobenius Manifold Structures on the Orbit Spaces of Affine Weyl Groups I

Abstract.
We present an approach to construct a class of generalized Frobenius manifold structures on the orbit spaces of affine Weyl groups, and prove that their monodromy groups are proper subgroups of the associated affine Weyl groups.
Key words and phrases:
Generalized Frobenius manifold, Affine Weyl group, Flat pencil of metrics, Monodromy Group
1. Introduction
The notion of Frobenius manifold was introduced by Dubrovin in the early 1990s in his study of the relation of integrable systems and 2-dimensional topological field theories [8, 9]. This geometric structure and its generalizations play a central role in topological field theory, Gromov-Witten theory, singularity theory and in the theory of integrable systems (see [1, 2, 3, 4, 5, 9, 13, 15, 18, 19, 22, 24] and the references therein).
We consider in this paper the construction of a class of generalized Frobenius manifold structures on the orbit spaces of affine Weyl groups. These structures satisfy all the axioms of Dubrovin’s definition of Frobenius manifolds except for the flatness of the unit vector field. The simplest example in this class corresponds to the affine Weyl group of type $A_{1}$, which naturally appears in the study of Hodge integrals and their remarkable connections to two important discrete integrable systems: the Volterra hierarchy and the $q$-deformed KdV hierarchy—both fundamental in soliton theory [21, 22].
Our approach parallels the construction of a Frobenius manifold structure on the orbit space of a Coxeter group or on a certain extension of an affine Weyl group [9, 12, 14, 25, 26, 28, 29]. For these cases, a flat pencil of metrics, based on which the Frobenius manifold structure is constructed, is obtained by projecting the Euclidean metric (on the space where the group acts) onto the orbit space. The key to these constructions lies in the linear dependence of both the contravariant components of the metric and its Levi-Civita connection on a distinguished coordinate of the orbit space—this special coordinate naturally gives rise to the flat unit vector field of the Frobenius manifold. However, for the generalized Frobenius manifolds that we are considering, the unit vector field is not flat, implying the absence of such a special coordinate. To overcome this, we proceed as follows: Given an affine Weyl group $W_{a}$ acting on an $\ell$-dimensional Euclidean space $V$, we introduce a parameter $\lambda$ into the invariant Fourier polynomials, yielding the coordinates $z^{1},\dots,z^{\ell}$ on the orbit space. The projected contravariant metric $\left(g^{\alpha\beta}(z)\right)$ then depends on $\lambda$. The key idea is to choose the $\lambda$-dependence of the invariant Fourier polynomials such that $\left(g^{\alpha\beta}(z)\right)$ varies linearly with $\lambda$.
In the present paper, we detail this construction, study properties of the resulting generalized Frobenius manifolds (including their monodromy groups), and provide some explicit examples. Further examples will appear in a subsequent paper [20]. The paper is organized as follows. In Sect. 2, we introduce the invariant $\lambda$-Fourier polynomial ring and study its properties. In Sect. 3, we construct generalized Frobenius manifold structures on orbit spaces of affine Weyl groups, and summarize the main results in Theorem 3.9. In Sect. 4, we analyze properties of the monodromy groups of these generalized Frobenius manifolds. In Sect. 5, we present some representative examples. In Sect. 6, we conclude the paper with remarks and further directions.
2. Invariant $\lambda$-Fourier polynomial rings and proper generators
Let us begin with several basic settings.
2.1. Affine coordinates
[…]
$\alpha_{j}^{\vee}=\frac{2\alpha_{j}}{(\alpha_{j},\alpha_{j})},\quad j=1,\dots,\ell,$
then the fundamental weights $\omega_{1},\dots,\omega_{\ell}\in V$ of $R$ are defined by the relations
$(\omega_{i},\alpha_{j}^{\vee})=\delta_{ij},\quad i,j=1,\dots,\ell.$
We fix a weight
$\omega=\sum_{j=1}^{\ell}m_{j}\omega_{j},\quad m_{j}\in\mathbb{Z}_{\geq 0}$ (2.1)
of $R$, and introduce affine coordinates $(x^{1},\dots,x^{\ell};c)$ on $V$ by
$\mathbf{x}=c\omega+x^{1}\alpha_{1}^{\vee}+\dots+x^{\ell}\alpha_{\ell}^{\vee},$ (2.2)
where $c\in\mathbb{R}$ is a fixed parameter. We will also use the following expressions of elements of $V$:
$\mathbf{x}=c\omega+\hat{\mathbf{x}}=(x^{1}+\theta_{1}c)\alpha_{1}^{\vee}+\dots+(x^{\ell}+\theta_{\ell}c)\alpha_{\ell}^{\vee},$ (2.3)
with $\hat{\mathbf{x}}=x^{1}\alpha_{1}^{\vee}+\dots+x^{\ell}\alpha_{\ell}^{\vee}$, and
$\theta_{j}=(\omega_{j},\omega),~{}j=1,\dots,\ell.$ (2.4)
Denote by $W_{a}(R)$ the affine Weyl group of $R$. It acts on $V$ by affine transformations
$\mathbf{x}\mapsto\sigma(\mathbf{x})+\sum_{j=1}^{\ell}k_{j}\alpha_{j}^{\vee},\quad\sigma\in W,\ k_{1},\dots,k_{\ell}\in\mathbb{Z},$ (2.5)
where $W=W(R)$ is the Weyl group of $R$ generated by the reflections
$\sigma_{j}(\mathbf{x})=\mathbf{x}-(\alpha_{j}^{\vee},\mathbf{x})\alpha_{j},\quad\mathbf{x}\in V,\ j=1,\dots,\ell.$
The action
$\sigma\left(c\omega+x^{1}\alpha_{1}^{\vee}+\dots+x^{\ell}\alpha_{\ell}^{\vee}\right)=c\omega+\tilde{x}^{1}\alpha_{1}^{\vee}+\dots+\tilde{x}^{\ell}\alpha_{\ell}^{\vee}$ (2.6)
of $\sigma\in W_{a}(R)$ on $V$ yields the change of affine coordinates
$\sigma(x^{1},\dots,x^{\ell};c)=(\tilde{x}^{1},\dots,\tilde{x}^{\ell};c).$
2.2. The Invariant $\lambda$-Fourier polynomial ring and its generators
Let us consider the Fourier function ring
$\mathscr{F}=\mathrm{span}_{\mathbb{C}}\left\{m(x^{1},\dots,x^{\ell};c)=\mathrm{e}^{2\pi i(t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell})}\mid t_{0},t_{1},\dots,t_{\ell}\in\mathbb{R}\right\}.$
We introduce a gradation on $\mathscr{F}$ by the assignment
$\deg\mathrm{e}^{2\pi ix^{j}}=\theta_{j},\quad\deg\mathrm{e}^{2\pi ic}=-1,$ (2.7)
where $\theta_{j}$ is defined in (2.4). Then for any $t_{0},t_{1},\dots,t_{\ell}\in\mathbb{R}$ and $\zeta\in\mathbb{C}$ we have
$\deg\mathrm{e}^{2\pi i(t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell}+\zeta)}=-t_{0}+t_{1}\theta_{1}+\dots+t_{\ell}\theta_{\ell}.$ (2.8)
In other words, for any given linear function $\psi(x^{1},\dots,x^{\ell};c)=t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell}$ and constant $m_{0}\in\mathbb{C}^{\times}$ we have
$\deg\left(m_{0}\mathrm{e}^{2\pi i\psi(x^{1},\dots,x^{\ell};c)}\right)=\psi(\theta_{1},\dots,\theta_{\ell};-1).$ (2.9)
The action of $W_{a}(R)$ on the affine coordinates of $V$ that is defined by (2.6) leads to a right action of $W_{a}(R)$ on $\mathscr{F}$, which is given by
$\sigma:~{}\mathscr{F}\to\mathscr{F},~{}f(x^{1},\dots,x^{\ell};c)\mapsto(f\circ\sigma)(x^{1},\dots,x^{\ell};c)$ (2.10)
for any given $\sigma\in W_{a}(R)$.
Lemma 2.1.
The affine Weyl group $W_{a}(R)$ acts homogeneously on $\mathscr{F}$, i.e., it preserves the degrees of monomials.
Proof.
Let $m(x^{1},\dots,x^{\ell};c)$ be a nonzero monimial, which can be represented in the form
$m(x^{1},\dots,x^{\ell};c)=m_{0}\mathrm{e}^{2\pi i\psi(x^{1},\dots,x^{\ell};c)}$
with a linear function $\psi(x^{1},\dots,x^{\ell};c)=t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell}$ and a nonzero constant $m_{0}$. For any given translation
$\sigma:\mathbf{x}\mapsto\mathbf{x}+\sum_{r=1}^{\ell}n_{r}\alpha_{r}^{\vee}$
of $W_{a}(R)$, its action on the affine coordinates $(x^{1},\dots,x^{\ell};c)$ is given by
$(x^{1},\dots,x^{\ell};c)\mapsto(x^{1}+n_{1},\dots,x^{\ell}+n_{\ell};c),$
so the associated action on $m_{0}\mathrm{e}^{2\pi i\psi(x^{1},\dots,x^{\ell};c)}$ yields $m_{0}^{\prime}\mathrm{e}^{2\pi i\psi(x^{1},\dots,x^{\ell};c)}$, which preserves the linear function $\psi(x^{1},\dots,x^{\ell};c)$ and only changes $m_{0}$ to another constant
$m_{0}^{\prime}=m_{0}\mathrm{e}^{2\pi i(t_{1}n_{1}+\dots+t_{\ell}n_{\ell})}.$
In particular, the action by $\sigma$ preserves the degree of $m(x^{1},\dots,x^{\ell};c)$.
On the other hand, the action of any given $\sigma\in W(R)$ on the affine coordinates $(x^{1},\dots,x^{\ell};c)$ is linear, hence
$(\sigma m)(x^{1},\dots,x^{\ell};c)=m_{0}\mathrm{e}^{2\pi i(\psi\circ\sigma)(x^{1},\dots,x^{\ell};c)}.$
From the relations
$\sigma(\mathbf{0})=\mathbf{0},\quad(-1)\omega+\theta_{1}\alpha_{1}^{\vee}+\dots+\theta_{\ell}\alpha_{\ell}^{\vee}=\mathbf{0},$
it follows that $\sigma$ preserves $(\theta_{1},\dots,\theta_{\ell};-1)$. Therefore by using (2.9) we arrive at
$\deg(\sigma m)=(\psi\circ\sigma)(\theta_{1},\dots,\theta_{\ell};-1)=\psi(\theta_{1},\dots,\theta_{\ell};-1)=\deg m.$
The lemma is proved. ∎
Denote by $\mathscr{F}^{W}$ the invariant subring of $\mathscr{F}$ w.r.t. the action of $W_{a}(R)$. Let
$Y_{j}(x^{1},\dots,x^{\ell};c)=\frac{1}{N_{j}}\sum_{\sigma\in W}\mathrm{e}^{2\pi i(\omega_{j},\sigma(\mathbf{x}))},\quad j=1,\dots,\ell,$ (2.11)
where $N_{j}=\#\{\sigma\in W\mid\sigma(\omega_{j})=\omega_{j}\}$ and $\mathbf{x}=c\omega+x^{1}\alpha_{1}^{\vee}+\dots+x^{\ell}\alpha_{\ell}^{\vee}$.
Proposition 2.2.
The functions $Y_{j}$ belong to $\mathscr{F}^{W}$ and are homogeneous of degree 00.
Proof.
The invariance of $Y_{j}$ under the action of $W_{a}(R)$ follows from their definition. For any $\sigma\in W_{a}(R)$, we know from Lemma 2.1 that
$\deg\mathrm{e}^{2\pi i(\omega_{j},\sigma(\mathbf{x}))}=\deg\mathrm{e}^{2\pi i(\omega_{j},\mathbf{x})}=\deg\mathrm{e}^{2\pi i(\theta_{j}c+x^{j})}=-\theta_{j}+\theta_{j}=0,$
thus we arrive at $\deg Y_{j}=0$. The proposition is proved. ∎
Let us proceed to consider generators of the $W_{a}(R)$-invariant ring $\mathscr{F}^{W}$. Since $\mathscr{F}^{W}$ is too large to consider its generators, we are to introduce a smaller subring of $\mathscr{F}$ which is named the $\lambda$-Fourier polynomial ring.
Denote
$\kappa=\gcd\{(\omega,\alpha_{r})~{}|~{}r=1,\dots,\ell\},$ (2.12)
where $\gcd$ refers to the greatest common divisor. This makes sense since
$(\omega,\alpha_{r})=m_{r}(\omega_{r},\alpha_{r})=m_{r}\frac{(\alpha_{r},\alpha_{r})}{2},\quad r=1,\dots,\ell$
are nonnegative rational numbers. Let
$\lambda=\mathrm{e}^{-2\pi i\kappa c},$ (2.13)
then we have
$\deg\lambda=\kappa.$ (2.14)
In particular, if we take $\omega=\omega_{k}$, then $\kappa=(\omega_{k},\alpha_{k})$ and $\lambda=\mathrm{e}^{-2\pi ic(\omega_{k},\alpha_{k})}$. Define the $\lambda$-Fourier polynomial ring $\mathscr{A}$ as a subring of $\mathscr{F}$ by
$\mathscr{A}=\mathbb{C}[\lambda]\otimes\mathbb{C}\bigl{[}\mathrm{e}^{2\pi ix^{j}},\mathrm{e}^{-2\pi ix^{j}}\mid j=1,\dots,\ell\bigr{]}\subset\mathscr{F}.$
Explicitly, $\mathscr{A}$ is composed of elements of the form
$\mathrm{e}^{2\pi i\left(-a_{0}\kappa c+a_{1}x^{1}+\dots+a_{\ell}x^{\ell}\right)},\quad a_{0}\in\mathbb{Z}_{\geq 0},a_{1},\dots,a_{\ell}\in\mathbb{Z}$
and their linear combinations. We denote by $\mathscr{A}^{W}=\mathscr{A}\cap\mathscr{F}^{W}$ the $W_{a}(R)$-invariant $\lambda$-Fourier polynomial ring. Then we have $\lambda=\mathrm{e}^{-2\pi i\kappa c}\in\mathscr{A}^{W}$.
Let
$y^{j}(\mathbf{x})=\mathrm{e}^{-2\pi i\theta_{j}c}~{}Y_{j}(\mathbf{x}),~{}j=1,\dots,\ell.$ (2.15)
From Proposition 2.2 it follows that $y^{j}$ is homogeneous of degree $\theta_{j}$.
Lemma 2.3.
We have $y^{j}\in\mathscr{A}^{W}$, $j=1,\dots,\ell$.
Proof.
It is obvious that $y^{j}$ is $W_{a}(R)$-invariant, so we only need to check that $y^{j}\in\mathscr{A}$. It suffices to show that for any $\sigma\in W(R)$, the monomial
$\mathrm{e}^{-2\pi i\theta_{j}c}\mathrm{e}^{2\pi i(\omega_{j},\sigma(\mathbf{x}))}\in\mathscr{A}.$ (2.16)
Observe that
$-2\pi i\theta_{j}c+2\pi i(\omega_{j},\sigma(\mathbf{x}))~{}=$ $~{}-2\pi ic(\omega_{j},\omega)+2\pi i(\sigma^{*}(\omega_{j}),\mathbf{x})$
$=$ $~{}-2\pi ic(\omega_{j}-\sigma^{*}(\omega_{j}),\omega)+2\pi i\sum_{r=1}^{\ell}x^{r}(\sigma^{*}(\omega_{j}),\alpha_{r}^{\vee}).$ (2.17)
Since $\sigma^{*}\in W(R)$, there exist non-negative integers $n_{1},\dots,n_{\ell}$ such that [6]
$\sigma^{*}(\omega_{j})=\omega_{j}-\sum_{s=1}^{\ell}n_{s}\alpha_{s},$ (2.18)
from which it follows that
$(\omega_{j}-\sigma^{*}(\omega_{j}),\omega)=\sum_{s=1}^{\ell}n_{s}\left(\alpha_{s},\omega\right)=m\kappa,$ (2.19)
where $m$ is a non-negative integer. We also have
$\left(\sigma^{*}(\omega_{j}),\alpha_{r}^{\vee}\right)=(\omega_{j},\alpha_{r}^{\vee})-\sum_{s=1}^{\ell}n_{s}(\alpha_{s},\alpha_{r}^{\vee})=\delta_{jr}-\sum_{s=1}^{\ell}n_{s}c_{sr},$ (2.20)
where $c_{sr}$ are the $(s,r)$-entries of the Cartan matrix of the root system $R$. Therefore, there exists $a_{0}\in\mathbb{Z}_{\geq 0}$ and $a_{1},\dots a_{\ell}\in\mathbb{Z}$ such that
$\mathrm{e}^{-2\pi i\theta_{j}c}\mathrm{e}^{2\pi i(\omega_{j},\sigma(\mathbf{x}))}=\mathrm{e}^{2\pi i(-a_{0}\kappa c+a_{1}x^{1}+\dots+a_{\ell}x^{\ell})}\in\mathscr{A}.$
The lemma is proved. ∎
Example 2.1 ($(R,\omega)=(A_{2},\omega_{2})$).
We consider the case where the root system $R=A_{2}$ and the weight $\omega=\omega_{2}$. We take the simple roots
$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3}$
of the root system $A_{2}$ in the hyperplane $V$ of the Euclidean space $\mathbb{R}^{3}$ spanned by $\alpha_{1},\alpha_{2}$. We have
$\alpha^{\vee}_{1}=\alpha_{1},\quad\alpha^{\vee}_{2}=\alpha_{2},\quad\omega_{1}=\frac{2}{3}\alpha^{\vee}_{1}+\frac{1}{3}\alpha^{\vee}_{2},\quad\omega_{2}=\frac{1}{3}\alpha^{\vee}_{1}+\frac{2}{3}\alpha^{\vee}_{2}.$
The affine coordinates $(x^{1},x^{2};c)$ on $V$ are given by
$\mathbf{x}=c\omega_{2}+x^{1}\alpha^{\vee}_{1}+x^{2}\alpha^{\vee}_{2},\quad\mathbf{x}\in V.$
The Weyl group $W$ acting on $V$ is generated by the reflections
$\sigma_{j}(\mathbf{x})=\mathbf{x}-(\alpha^{\vee}_{j},\mathbf{x})\alpha_{j},\quad j=1,2,\ \mathbf{x}\in V.$
From the definition (2.11) of $Y_{j}$, we have
$Y_{1}$ $=\mathrm{e}^{2\pi i\left(x^{1}+\frac{1}{3}c\right)}+\mathrm{e}^{2\pi i\left(x^{2}-x^{1}+\frac{1}{3}c\right)}+\mathrm{e}^{-2\pi i\left(x^{2}+\frac{2}{3}c\right)},$
$Y_{2}$ $=\mathrm{e}^{2\pi i\left(x^{2}+\frac{2}{3}c\right)}+\mathrm{e}^{-2\pi i\left(x^{1}+\frac{1}{3}c\right)}+\mathrm{e}^{2\pi i\left(x^{1}-x^{2}-\frac{1}{3}c\right)}.$
Here $\kappa=1$, hence
$\lambda=\mathrm{e}^{-2\pi ic}.$
Then we have
$y^{1}$ $=e^{-\frac{2}{3}\pi ic}Y_{1}=\mathrm{e}^{2\pi ix^{1}}+\mathrm{e}^{2\pi i(x^{2}-x^{1})}+\lambda\mathrm{e}^{-2\pi ix^{2}},$
$y^{2}$ $=e^{-\frac{4}{3}\pi ic}Y_{2}=\mathrm{e}^{2\pi ix^{2}}+\lambda\mathrm{e}^{-2\pi ix^{1}}+\lambda\mathrm{e}^{2\pi i(x^{1}-x^{2})}.$
In the next subsection we are to show that $y^{1},\dots,y^{\ell}$ and $\lambda$ generate $\mathscr{A}^{W}$.
2.3. The $\varphi_{\lambda}$-transformation and the null transformation
The parameter $\lambda$ of the $\lambda$-Fourier polynomials $y^{1},\dots,y^{\ell}$ is defined by (2.13), so it lies on the unit circle. However, since $y^{j}$ depend polynomially on $\lambda$, the definition of $y^{j}$ can be extended naturally to any $\lambda\in\mathbb{C}$, including $\lambda=0$.
Definition 2.1.
For any fixed $\lambda\in\mathbb{C}$, the $\lambda$-Fourier polynomials $y^{1},\dots,y^{\ell}$ yield a transformation
$\varphi_{\lambda}\colon(x^{1},\dots,x^{\ell})\mapsto(y^{1},\dots,y^{\ell}).$
We call it the $\varphi_{\lambda}$-transformation of $(y^{1},\dots,y^{\ell})$. In particular, the $\varphi_{0}$ transformation is called the null transformation.
Example 2.2 ($(R,\omega)=(A_{2},\omega_{2})$).
From the Example 2.1, we know that the $\varphi_{1}$-transformation of $(y^{1},y^{2})$ is given by
$y^{1}\big{|}_{\lambda=1}$ $=\mathrm{e}^{2\pi ix^{1}}+\mathrm{e}^{2\pi i(x^{2}-x^{1})}+\mathrm{e}^{-2\pi ix^{2}},$
$y^{2}\big{|}_{\lambda=1}$ $=\mathrm{e}^{2\pi ix_{2}}+\mathrm{e}^{-2\pi ix^{1}}+\mathrm{e}^{2\pi i(x^{1}-x^{2})}.$
The corresponding null transformation is given by
$y^{1}\big{|}_{\lambda=0}$ $=\mathrm{e}^{2\pi ix^{1}}+\mathrm{e}^{2\pi i(x^{2}-x^{1})},$
$y^{2}\big{|}_{\lambda=0}$ $=\mathrm{e}^{2\pi ix^{2}}.$
Note that $y^{2}$ has only one term $\mathrm{e}^{2\pi ix^{2}}$, this result will be generalized in Theorem 2.7.
Proposition 2.4.
The Jacobian determinant of the null transformation is given by
$\mathbf{J}_{0}(\hat{\mathbf{x}})=e^{\pi i\sum_{\alpha\in\Phi^{+}}(\alpha,\hat{\mathbf{x}})}\prod_{\beta\in\Phi^{+},\,(\beta,\omega)=0}\left(1-\mathrm{e}^{-2\pi i(\beta,\hat{\mathbf{x}})}\right),$ (2.21)
where $\Phi^{+}$ is the set of all positive roots of $R$.
Proof.
From [6] and [12] we know that the Jacobian determinant of $\varphi_{\lambda}$ can be represented in the form
$\mathbf{J}_{\lambda}(\hat{\mathbf{x}})~{}=$ $~{}\mathrm{e}^{-2\pi ic\sum_{j}(\omega_{j},\omega)+\pi i\sum_{\alpha\in\Phi^{+}}\left(c(\alpha,\omega)+(\alpha,\hat{\mathbf{x}})\right)}\prod_{\beta\in\Phi^{+}}\left(1-\mathrm{e}^{-2\pi i\left(c(\beta,\omega)+(\beta,\hat{\mathbf{x}})\right)}\right)$
$=$ $~{}e^{\pi i\sum_{\alpha\in\Phi^{+}}(\alpha,\hat{\mathbf{x}})}\prod_{\beta\in\Phi^{+}}\left(1-\mathrm{e}^{-2\pi i\left(c(\beta,\omega)+(\beta,\hat{\mathbf{x}})\right)}\right).$ (2.22)
Here we use the fact that [6]
$\frac{1}{2}\sum_{\alpha\in\Phi^{+}}\alpha=\sum_{j=1}^{\ell}\omega_{j}.$ (2.23)
Since $\mathbf{J}_{\lambda}(\hat{\mathbf{x}})$ is a polynomial of $\lambda$, its constant term w.r.t. $\lambda$ yields (2.21). The proposition is proved. ∎
Corollary 2.5.
The Jacobian determinant of the null transformation is non-vanishing except on an $(\ell-1)$-dimensional divisor. In particular, $y^{1}\big{|}_{\lambda=0},\dots,y^{\ell}\big{|}_{\lambda=0}$ are functionally independent.
Theorem 2.6.
We have $\mathscr{A}^{W}=\mathbb{C}[y^{1},\dots,y^{\ell};\lambda]$.
Proof.
For any given $f\in\mathscr{A}^{W}$, we need to show that it can be represented in terms of a polynomial of $y^{1},\dots,y^{\ell}$ and $\lambda$. By using Lemma 2.1, we know that $f$ can be decomposed as the sum of homogeneous elements of $\mathscr{A}^{W}$, so we may assume without lose of generality that $f$ is homogeneous of degree $n$.
From [6] we know that the Fourier polynomial $f|_{\lambda=1}$ of $x^{1},\dots,x^{\ell}$ can be represented in terms of a polynomial of $y^{1}\big{|}_{\lambda=1}$,…,$y^{\ell}\big{|}_{\lambda=1}$ as follows:
$f|_{\lambda=1}=\sum_{\beta_{1},\dots,\beta_{\ell}\geq 0}a_{\beta_{1},\dots,\beta_{\ell}}(y^{1}\big{|}_{\lambda=1})^{\beta_{1}}\dots(y^{\ell}\big{|}_{\lambda=1})^{\beta_{\ell}},$ (2.24)
so it follows from the homogeneity of $f$ and of $y^{1},\dots,y^{\ell}$ that
$f=\sum_{\beta_{1},\dots,\beta_{\ell}\geq 0}a_{\beta_{1},\dots,\beta_{\ell}}\lambda^{n-\theta_{1}\beta_{1}-\dots-\theta_{\ell}\beta_{\ell}}(y^{1})^{\beta_{1}}\cdots(y^{\ell})^{\beta_{\ell}},$ (2.25)
here $n-\theta_{1}\beta_{1}-\dots-\theta_{\ell}\beta_{\ell}\in\mathbb{Z}$. Hence we can represent $f$ in the form
$f=\sum_{j\in Z}\lambda^{j}P_{j}(y^{1},\dots,y^{\ell}),$ (2.26)
where $P_{j}$ are polynomials. Let $q$ be the smallest integer such that $P_{q}=0$. If $q<0$, then by viewing $f$ and $y^{1},\dots,y^{\ell}$ as functions of $x^{1},\dots,x^{\ell}$ and $\lambda$, and by comparing the $\lambda^{q}$ term in (2.26) we arrive at
$P_{q}(y^{1}\big{|}_{\lambda=0},\dots,y^{\ell}\big{|}_{\lambda=0})=0.$ (2.27)
This leads to a contradiction with the fact, as shown by Corollary 2.5, that $y^{1}\big{|}_{\lambda=0},\dots,y^{\ell}\big{|}_{\lambda=0}$ are functionally independent. Thus $q$ must be a non-negative integer, and the proposition is proved. ∎
Definition 2.2.
We call $\{y^{1},\dots,y^{\ell}\}$ the set of basic generators of $\mathscr{A}^{W}$.
2.4. Leading terms of the basic generators
From the definition (2.15) of $y^{j}$, we know that it contains a term
$\mathrm{e}^{-2\pi i\theta_{j}c}\mathrm{e}^{2\pi i(\omega_{j},{\mathbf{x}})}=\mathrm{e}^{2\pi ix^{j}},$ (2.28)
which we call the leading term of $y^{j}$. This term also appears in $y^{j}|_{\lambda=0}$, and the following theorem shows that for some indices $j$, $y^{j}|_{\lambda=0}$ only contain their leading terms.
Theorem 2.7.
Let us denote
$S=\{r\in\{1,\dots,\ell\}\mid m_{r}=(\omega,\alpha_{r}^{\vee})>0\}.$ (2.29)
Then we have
$y^{j}\big{|}_{\lambda=0}=\mathrm{e}^{2\pi ix^{j}},\quad j\in S.$ (2.30)
Proof.
For any given $j\in S$, from (2.11) and (2.15) we obtain
$y^{j}\big{|}_{\lambda=0}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W\\ (\omega_{j}-\sigma^{*}(\omega_{j}),\omega)=0\end{subarray}}\mathrm{e}^{2\pi i(\sigma^{*}(\omega_{j}),\hat{\mathbf{x}})}.$ (2.31)
Assume that $\sigma\in W$ satisfies the condition
$(\omega_{j}-\sigma^{*}(\omega_{j}),\omega)=0,$ (2.32)
then from (2.18) it follows that
$\sum_{r=1}^{\ell}n_{r}m_{r}(\alpha_{r},\omega_{r})=0.$
By using the fact that all terms on the left hand side of the above equality are nonnegative and $m_{j}>0$, we obtain $n_{j}=0$, and
$(\sigma^{*}(\omega_{j}),\omega_{j})=(\omega_{j},\omega_{j})-\sum_{r=1}^{\ell}n_{r}(\alpha_{r},\omega_{j})=(\omega_{j},\omega_{j}).$ (2.33)
Since $\sigma^{*}$ is an orthogonal transformation, the equation (2.33) implies that $\sigma^{*}(\omega_{j})=\omega_{j}$. Therefore, from (2.31) it follows that
$y^{j}\big{|}_{\lambda=0}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W\\ \omega_{j}=\sigma^{*}(\omega_{j})\end{subarray}}\mathrm{e}^{2\pi i(\omega_{j},\hat{\mathbf{x}})}=\frac{1}{N_{j}}\sum_{\begin{subarray}{c}\sigma\in W\\ \omega_{j}=\sigma^{*}(\omega_{j})\end{subarray}}\mathrm{e}^{2\pi ix^{j}}=\mathrm{e}^{2\pi ix^{j}}.$ (2.34)
The theorem is proved. ∎
Corollary 2.8.
The following identity holds true:
$\sum_{r=1}^{\ell}m_{r}\log\left(y^{r}\big{|}_{\lambda=0}\right)=2\pi i(\omega,\hat{\mathbf{x}}).$ (2.35)
2.5. Proper generators
In order to construct generalized Frobenius manifold structures on the orbit space of $W_{a}(R)$, we introduce in this subsection some other generators of $\mathscr{A}^{W}$.
Definition 2.3.
Suppose $z^{1},\dots,z^{\ell}\in\mathscr{A}^{W}$. We say that $\{z^{1},\dots,z^{\ell}\}$ is a set of proper generators of $\mathscr{A}^{W}$ if $\deg z^{j}=\theta_{j}$ and
$z^{j}\big{|}_{\lambda=0}=y^{j}\big{|}_{\lambda=0},\quad j=1,\dots,\ell.$ (2.36)
Lemma 2.9.
Let $z^{1},\dots,z^{\ell}\in\mathscr{A}^{W}$, then these functions form a set of proper generators of $\mathscr{A}^{W}$ if and only if for any $1\leq j\leq\ell$ there exists a polynomial $s^{j}$ of $\lambda$ and $\{y^{r}\big{|}~{}1\leq r\leq\ell,~{}\theta_{r}<\theta_{j}\}$ such that either $s^{j}=0$ or $\deg s^{j}=\theta_{j}-\kappa$, and
$z^{j}=y^{j}+\lambda s^{j}.$ (2.37)
Proof.
If $z^{1},\dots,z^{\ell}$ satisfy the relation (2.37), then it is obvious to see that they form a set of proper generators of $\mathscr{A}^{W}$.
Now we suppose that $\{z^{1},\dots,z^{\ell}\}$ is a set of proper generators of $\mathscr{A}^{W}$, then for any $1\leq j\leq\ell$ we have $(z^{j}-y^{j})\in\mathscr{A}^{W}$ and
$\left(z^{j}-y^{j}\right)\big{|}_{\lambda=0}=0.$
Hence there exists $s^{j}\in\mathscr{A}^{W}$ such that $z^{j}-y^{j}=\lambda s^{j}$, and either $s^{j}=0$ or $\deg s^{j}=\theta_{j}-\kappa$. By using Theorem 2.6 and by degree counting, we know that $s^{j}$ is a polynomial of $\lambda$ and $y^{r}$’s whose degrees are less than $\theta_{j}$. The lemma is proved. ∎
Proposition 2.10.
Let $\{z^{1},\dots,z^{\ell}\}$ be a set of proper generators of $\mathscr{A}^{W}$, then
$\mathscr{A}^{W}=\mathbb{C}[z^{1},\dots,z^{\ell},\lambda].$ (2.38)
Proof.
It suffices to show that $y^{j}\in\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$ for $j=1,\dots,\ell$. Assume that there exist some $y^{j}$ which do not belong to $\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$, and let $y^{k}$ be one of those with the lowest degree. It follows from (2.37) that $s^{k}\not\in\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$. However, since $\theta_{k}$ is the lowest degree among the ones of $y^{j}\not\in C[z^{1},\dots,z^{\ell},\lambda]$, we know that $y^{r}\in\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$ for each $r$ satisfying $\theta_{r}<\theta_{k}$. Then we know from Lemma 2.9 that $s^{k}\in\mathbb{C}[z^{1},\dots,z^{\ell},\lambda]$, which leads to a contradiction. The proposition is proved. ∎
Corollary 2.11.
For any set $\{z^{1},\dots,z^{\ell}\}$ of proper generators of $\mathscr{A}^{W}$, we have
$\det\left(\frac{\partial z^{j}}{\partial x^{p}}\right)=\det\left(\frac{\partial y^{j}}{\partial x^{p}}\right)=\mathbf{J}_{\lambda}(\hat{\mathbf{x}}),$ (2.39)
where $\mathbf{J}_{\lambda}(\hat{\mathbf{x}})$ is defined in (2.22), thus $z^{1},\dots,z^{\ell}$ are functionally independent for each $\lambda\in\mathbb{C}$.
Proof.
For each $1\leq j\leq\ell$ we have $z^{j}=y^{j}+\lambda s^{j}$, where $s^{j}$ is a certain polynomial of $\lambda$ and $y^{r}$ satisfying the condition $\theta_{r}<\theta_{j}$. Then from the relation
$\frac{\partial z^{j}}{\partial x^{p}}=\frac{\partial y^{j}}{\partial x^{p}}+\lambda\sum_{\theta_{r}<\theta_{j}}\frac{\partial s^{j}}{\partial y^{r}}\frac{\partial y^{r}}{\partial x^{p}}$ (2.40)
we arrive at (2.39). The corollary is proved. ∎
For any given set $\{z^{1},\dots,z^{\ell}\}$ of proper generators of $\mathscr{A}^{W}$, we will also call the transformation which maps $(x^{1},\dots,x^{\ell})$ to $(z^{1},\dots,z^{\ell})$ the $\varphi_{\lambda}$-transformation, and the corresponding $\varphi_{0}$-transformation the null transformation. Note that $\{z^{1},\dots,z^{\ell}\}$ and $\{y^{1},\dots,y^{\ell}\}$ may lead to distinct $\varphi_{\lambda}$-transformations, but they always share the same null-transformation, i.e.,
$z^{j}\big{|}_{\lambda=0}=y^{j}\big{|}_{\lambda=0},\quad j=1,\dots,\ell.$ (2.41)
In particular, for any element $r$ of the set $S$ defined by (2.29) we have
$z^{r}\big{|}_{\lambda=0}=\mathrm{e}^{2\pi ix^{r}}.$ (2.42)
Hence from Corollary 2.35 it follows that
$\sum_{r=1}^{\ell}m_{r}\log\left(z^{r}\big{|}_{\lambda=0}\right)=2\pi i(\omega,\hat{\mathbf{x}}).$ (2.43)
In the next section, we will impose an additional condition on proper generators of $\mathscr{A}^{W}$ in order to construct generalized Frobenius manifold structures on the orbits space of the affine Weyl group $W_{a}(R)$.
3. The construction of generalized Frobenius manifold structures
In this section, we are to construct a generalized Frobenius manifold structure starting from a given irreducible reduced root system $R$ in the $l$-dimensional Euclidean space $V$, a fixed weight $\omega$ of the form (2.1), and a set $\{z^{1},\dots,z^{\ell}\}$ of proper generators of $\mathscr{A}^{W}$ satisfying a certain additional condition.
3.1. The orbit space and pencil generators
By definition, $\mathbb{C}[z^{1},\cdots,z^{\ell}]$ is a subalgebra of $\mathscr{A}^{W}$. Denote the spectrum of this subalgebra by $\mathcal{M}$, then $\mathcal{M}\cong\mathbb{C}^{\ell}$. Taking $x^{1},\cdots,x^{\ell}$ as coordinates on the complexified Euclidean space $V\otimes\mathbb{C}$, we have a collection of holomorphic maps
$\varphi_{\lambda}\colon V\otimes\mathbb{C}\to\mathcal{M},\quad(x^{1},\dots,x^{\ell})\mapsto(z^{1},\dots,z^{\ell}),\quad\lambda\in\mathbb{C}.$ (3.1)
Here the $\varphi_{\lambda}$-transformation is defined in the last section. We call $\mathcal{M}$ the orbit space of the affine Weyl group $W_{a}(R)$ w.r.t. the proper generators $z^{1},\dots,z^{\ell}$. In this section, we will use this collection of maps $\{\varphi_{\lambda}\}_{\lambda\in\mathbb{C}}$ to endow $\mathcal{M}$ with a structure of generalized Frobenius manifold.
[…]
$a(\mathrm{d}x^{i},\mathrm{d}x^{j})=a^{ij}\ \mbox{with}\ (a^{ij})=\left((\alpha_{i}^{\vee},\alpha_{j}^{\vee})\right)^{-1},\quad i,j=1,\dots,\ell$ (3.2)
on $V\otimes\mathbb{C}$. Here by a contravariant metric we mean a symmetric non-degenerate bilinear form on the holomorphic cotangent bundle $T^{*}(V\otimes\mathbb{C})$.
Proposition 3.1.
The numbers $\theta_{j}$ defined by (2.4) can be represented in the form
$\theta_{j}=a^{jr}m_{r},\quad j=1,\dots,\ell.$ (3.3)
Here and in what follows, we assume summation over repeated upper and lower indices.
Proof.
From $\omega=\sum_{s}m_{s}\omega_{s}=\sum_{p}\theta_{p}\alpha_{p}^{\vee}$ it follows that
$\sum_{p=1}^{\ell}\theta_{p}(\alpha_{p}^{\vee},\alpha_{s}^{\vee})=(\omega,\alpha_{s}^{\vee})=m_{s},$ (3.4)
thus (3.3) holds true. The proposition is proved. ∎
The $\varphi_{\lambda}$-transformation induces a collection of pushing forward $\{(\varphi_{\lambda})_{*}\}_{\lambda\in\mathbb{C}}$, which transforms $(a^{ij})$ to a contravariant metric $(\varphi_{\lambda})_{*}a$ on $\mathcal{M}$. We denote
$g_{\lambda}=\frac{1}{4\pi^{2}}(\varphi_{\lambda})_{*}a=\bigl{(}g_{\lambda}^{ij}\bigr{)},$
where
$g_{\lambda}^{ij}=\frac{1}{4\pi^{2}}\frac{\partial z^{i}}{\partial x^{r}}a^{rs}\frac{\partial z^{j}}{\partial x^{s}},\quad i,j=1,\dots,\ell.$ (3.5)
Note that we introduce the factor $1/{4\pi^{2}}$ to simplify the expressions of $g_{\lambda}^{ij}$’s.
Proposition 3.2.
We have $g^{ij}_{\lambda}\in\mathscr{A}^{W}$.
Proof.
By definition, we have $g^{ij}_{\lambda}\in\mathscr{A}$, hence it suffices to show that $g^{ij}_{\lambda}$ is invariant w.r.t. the action of $W_{a}(R)$. Since the invariance of $g^{ij}_{\lambda}$ w.r.t. translations is obvious, we consider actions by elements of the Weyl group $W(R)$. Write $z^{j}$ as $z^{j}=z^{j}(x^{1},\dots,x^{\ell};c)$. For any given $\sigma\in W(R)$, since $z^{j}(x^{1},\dots,x^{\ell};c)=z^{j}(\sigma(x^{1},\dots,x^{\ell};c))$, we have
$\frac{\partial z^{j}}{\partial x^{r}}(x^{1},\dots,x^{\ell};c)=\frac{\partial z^{j}}{\partial x^{p}}(\sigma(x^{1},\dots,x^{\ell};c))\frac{\partial\tilde{x}^{p}}{\partial x^{r}},$ (3.6)
where $\sigma(x^{1},\dots,x^{\ell};c)=(\tilde{x}^{1},\dots,\tilde{x}^{\ell};c)$. Note that $\frac{\partial\tilde{x}^{p}}{\partial x^{r}}$ are constants since $\sigma$ is a linear transformation. Therefore,
$g^{ij}_{\lambda}(x^{1},\dots,x^{\ell};c)=\frac{1}{4\pi^{2}}\frac{\partial z^{i}}{\partial x^{r}}(x^{1},\dots,x^{\ell};c)a^{rs}\frac{\partial z^{j}}{\partial x^{s}}(x^{1},\dots,x^{\ell};c)$
$\quad=\frac{1}{4\pi^{2}}\frac{\partial z^{i}}{\partial x^{p}}(\sigma(x^{1},\dots,x^{\ell};c))\frac{\partial z^{j}}{\partial x^{q}}(\sigma(x^{1},\dots,x^{\ell};c))\frac{\partial\tilde{x}^{p}}{\partial x^{r}}a^{rs}\frac{\partial\tilde{x}^{q}}{\partial x^{s}}.$
Since $\sigma$ is an orthogonal transformation, we know that
$\frac{\partial\tilde{x}^{p}}{\partial x^{r}}a^{rs}\frac{\partial\tilde{x}^{q}}{\partial x^{s}}=a^{pq}.$
Hence $g^{ij}_{\lambda}(x^{1},\dots,x^{\ell};c)=g^{ij}_{\lambda}(\sigma(x^{1},\dots,x^{\ell};c))$. The proposition is proved. ∎
The above proposition shows in particular that $g^{ij}_{\lambda}$ are polynomials of $\lambda$. In order to construct generalized Frobenius manifold structures on the orbit space $\mathcal{M}$, we are to impose an additional condition on the proper generators of $\mathscr{A}^{W}$, so that $g_{\lambda}$ yields a flat pencil of metrics [10] on $\mathcal{M}$.
Definition 3.1.
Suppose $\{z^{1},\dots,z^{\ell}\}$ be a set of proper generators of $\mathscr{A}^{W}$. We say that it is a set of pencil generators of $\mathscr{A}^{W}$ if the associated metric $g_{\lambda}=\bigl{(}g^{ij}_{\lambda}\bigr{)}$ depends linearly on $\lambda$, i.e., $g^{ij}_{\lambda}$ can be represented in the form
$g^{ij}_{\lambda}=g^{ij}+\lambda\eta^{ij},$
where $\eta=(\eta^{ij})$ is non-degenerate at generic points of $\mathcal{M}$, and the contravariant components $\Gamma^{ij}_{g_{\lambda},k}=-g_{\lambda}^{ir}\Gamma_{g_{\lambda},rk}^{j}$ of the Levi-Civita connection of $g_{\lambda}$ can be represented in terms of that of $g$ and $\eta$ as follows:
$\Gamma^{ij}_{g_{\lambda},k}=\Gamma^{ij}_{g,k}+\lambda\Gamma^{ij}_{\eta,k}.$ (3.7)
From the above definition we know that a set of pencil generators of $\mathscr{A}^{W}$ yields a flat pencil[10] of metrics $g,\eta$. Note that the flat metric $g$ is independent of the choice of proper generators, since
$g=g_{0}=\frac{1}{4\pi^{2}}(\varphi_{0})_{*}a$ (3.8)
is induced merely by the null-transformation which does not depend on the choice of proper generators.
We will present examples of pencil generators in Sect. 5 and in a subsequent paper [20]. Unless otherwise specified, in what follows we always assume that $\{z^{1},\dots,z^{\ell}\}$ is a set of pencil generators of $\mathscr{A}^{W}$. We will end this section by proving that we can construct a generalized Frobenius manifold structure on $\mathcal{M}$ with the given flat pencil of metrics.
3.2. Homogeneous flat coordinates
We are to show in this subsection the existence of a system of homogeneous flat coordinates for the flat metric $\eta$.
Flat coordinates of $\eta$ satisfy the system of equations
$\frac{\partial^{2}t}{\partial z^{i}\partial z^{j}}=\Gamma_{\eta,ij}^{k}\frac{\partial t}{\partial z^{k}},\quad i,j=1,\dots,\ell.$ (3.9)
Here $\Gamma_{\eta,ij}^{k}$ are the Christoffel symbols of $\eta$’s Levi-Cività connection, they are given by
$\Gamma_{\eta,ij}^{k}=\frac{1}{2}\eta^{kr}\left(\frac{\partial\eta_{ir}}{\partial z^{j}}+\frac{\partial\eta_{jr}}{\partial z^{i}}-\frac{\partial\eta_{ij}}{\partial z^{r}}\right).$ (3.10)
From Definition 2.3 we know that
$\deg g^{ij}=\theta_{i}+\theta_{j},\quad\deg\eta^{ij}=\theta_{i}+\theta_{j}-\kappa,$ (3.11)
which lead to
$\deg g_{ij}=-\theta_{i}-\theta_{j},\quad\deg\eta_{ij}=\kappa-\theta_{i}-\theta_{j},\quad\deg\Gamma_{\eta,ij}^{k}=\theta_{k}-\theta_{i}-\theta_{j}.$ (3.12)
We are to use the above homogeneity property to prove the existence of homogeneous flat coordinates of $\eta$. To this end, let us first clarify the solution space of the system of equations (3.9) that we are considering. Let $\widehat{\mathcal{M}}$ be the universal covering of $\mathcal{M}\setminus D$ with
$D=\{z\in\mathcal{M}~{}\big{|}~{}\det\eta(z)=0\}.$ (3.13)
We consider solutions of the system of equations (3.9) on the space
$\mathscr{Q}(\widehat{\mathcal{M}})=\mathscr{O}(\widehat{\mathcal{M}})/\mathbb{C},$
where $\mathscr{O}(\widehat{\mathcal{M}})$ refers to the holomorphic function ring on $\mathcal{M}$. Denote the solution space of (3.9) by
$\mathcal{K}=\left\{t\in\mathscr{Q}(\widehat{\mathcal{M}})\mid\frac{\partial^{2}t}{\partial z^{i}\partial z^{j}}=\Gamma_{\eta,ij}^{k}\frac{\partial t}{\partial z^{k}},\ i,j=1,\dots,\ell\right\}.$
It follows from the theory of differential equations that $\mathcal{K}$ is an $\ell$-dimensional space, since a solution is uniquely determined by the values of its $\ell$ partial derivatives at an arbitrary point.
Now we introduce the Euler vector field
$E=\sum_{r=1}^{\ell}\theta_{r}z^{r}\frac{\partial}{\partial z^{r}}.$ (3.14)
Then for any homogeneous function $f$ of degree $n$, the following identity holds true:
$\mathcal{L}_{E}f=nf,$ (3.15)
here $\mathcal{L}_{E}$ is the Lie derivative along $E$. Hence homogeneous functions can be seen as eigenvectors of $\mathcal{L}_{E}$, of which the eigenvalues are their degrees. Now we consider the action of $\mathcal{L}_{E}$ on the solution space $\mathcal{K}$.
Proposition 3.3.
The space $\mathcal{K}$ is $\mathcal{L}_{E}$-invariant, i.e., $\mathcal{L}_{E}(\mathcal{K})\subseteq\mathcal{K}$.
Proof.
Denote
$\mathcal{L}_{ij}=\frac{\partial^{2}}{\partial z^{i}\partial z^{j}}-\Gamma_{\eta,ij}^{k}\frac{\partial}{\partial z^{k}}.$ (3.16)
By using (3.12) we obtain
$\mathcal{L}_{E}\Gamma_{\eta,ij}^{k}=(\theta_{k}-\theta_{i}-\theta_{j})\Gamma_{\eta,ij}^{k},$
from which it follows that
$\left[\mathcal{L}_{E},\mathcal{L}_{ij}\right]=-(\theta_{i}+\theta_{j})\mathcal{L}_{ij}.$ (3.17)
Hence if $t\in\ker\mathcal{L}_{ij}$, then $\mathcal{L}_{E}t\in\ker\mathcal{L}_{ij}$. The proposition is proved. ∎
In what follows, we always assume that $\mathcal{L}_{E}|_{\mathcal{K}}$ is diagonalizable. This implies that $\eta$ admits $\ell$ homogeneous flat coordinates which we denote by $t^{1},\dots,t^{\ell}$. Let
$d_{\alpha}=\deg t^{\alpha},\quad\alpha=1,\dots,\ell,$ (3.18)
then we have
$E=\sum_{r=1}^{\ell}\theta_{r}z^{r}\frac{\partial t^{\alpha}}{\partial z^{r}}\frac{\partial}{\partial t^{\alpha}}=\sum_{\alpha=1}^{\ell}d_{\alpha}t^{\alpha}\frac{\partial}{\partial t^{\alpha}}.$ (3.19)
Now let us consider the flat pencil of metrics $g$ and $\eta$ in the flat coordinates $t^{1},\dots,t^{\ell}$. We denote their components by $g^{\alpha\beta}$ and $\eta^{\alpha\beta}$ with Greek indices respectively. Recall that we use $g^{ij}$ and $\eta^{ij}$ with Latin indices to denote the components of $g$ and $\eta$ in the coordinates $z^{1},\dots,z^{\ell}$. Since $\eta^{\alpha\beta}$ are constants, from
$\deg\eta^{\alpha\beta}=d_{\alpha}+d_{\beta}-\kappa$ (3.20)
it follows that if $d_{\alpha}+d_{\beta}\not=\kappa$ then $\eta^{\alpha\beta}$ must vanish. Thus by performing, if necessary, a degree preserving linear transformation, we can choose the flat coordinates $t^{1},\dots,t^{n}$ of $\eta$ such that
$\eta^{\alpha\beta}=\delta_{\beta,\alpha^{*}}.$ (3.21)
Here * is an involution
:$*\colon\{1,\dots,l\}\to\{1,\dots,l\},\quad\alpha\to\alpha^{*},$
which satisfies the relation $d_{\alpha}+d_{\alpha^{*}}=\kappa$. Usually, we prefer to choose the flat coordinates so that $\alpha^{*}=\ell+1-\alpha$.
We note that $d_{1},\dots,d_{\ell}$ are the eigenvalues of the operator $\nabla^{\eta}E$ corresponding to the eigenvectors $\frac{\partial}{\partial t^{1}},\dots,\frac{\partial}{\partial t^{\ell}}$, so they are also the eigenvalues of the representation matrix $A$ of $\nabla^{\eta}E$ in the basis $\frac{\partial}{\partial z^{1}},\dots,\frac{\partial}{\partial z^{\ell}}$. Since
$\nabla_{\frac{\partial}{\partial z^{i}}}^{\eta}E=\nabla_{\frac{\partial}{\partial z^{i}}}^{\eta}\sum_{k=1}^{\ell}\theta_{k}z^{k}\frac{\partial}{\partial z^{k}}=\left(\sum_{k=1}^{\ell}\theta_{k}z^{k}\Gamma_{\eta,ik}^{j}+\theta_{i}\delta^{j}_{i}\right)\frac{\partial}{\partial z^{j}},$ (3.22)
the matrix $A$ has elements
$A_{j}^{i}=\sum_{k=1}^{\ell}\theta_{k}z^{k}\Gamma_{\eta,jk}^{i}+\theta_{j}\delta^{i}_{j}.$ (3.23)
It has eigenvectors $\left(\frac{\partial t^{\alpha}}{\partial z^{1}},\dots,\frac{\partial t^{\alpha}}{\partial z^{\ell}}\right)^{T}$ with eigenvalues $d_{\alpha}$, i.e.,
$A\begin{pmatrix}\frac{\partial t^{\alpha}}{\partial z^{1}}\\ \vdots\\ \frac{\partial t^{\alpha}}{\partial z^{r}}\end{pmatrix}=d_{\alpha}\begin{pmatrix}\frac{\partial t^{\alpha}}{\partial z^{1}}\\ \vdots\\ \frac{\partial t^{\alpha}}{\partial z^{r}}\end{pmatrix}.$ (3.24)
Remark 3.1.
When we choose $\omega$ to be a fundamental weight, we conjecture that all $d_{\alpha}$ are positive numbers. In what follows, we will continue our construction by assuming this fact.
3.3. The Frobenius algebra
Denote by $\Gamma_{\alpha\gamma}^{\beta}$ the Christoffel symbols of the Levi-Civita connection of $g$ in the flat coordinates $t^{1},\dots,t^{\ell}$ of $\eta$, and by
$\Gamma_{\gamma}^{\alpha\beta}=-g^{\alpha\epsilon}\Gamma_{\epsilon\gamma}^{\beta}$ (3.25)
the contravariant components of the Levi-Civita connection of $g$. Then from (3.7) we know that $\Gamma_{\gamma}^{\alpha\beta}$ are also the contravariant components of the Levi-Civita connection of $g+\lambda\eta$ for each $\lambda\in\mathbb{C}$.
We define an algebra structure on $T^{*}\mathcal{M}$ by
$\mathrm{d}t^{\alpha}\cdot\mathrm{d}t^{\beta}=\frac{\kappa}{d_{\beta}}\Gamma_{\gamma}^{\alpha\beta}\mathrm{d}t^{\gamma},\quad\alpha,\beta=1,\dots,\ell,$ (3.26)
[…]
$\omega_{e}=-\sum_{r=1}^{\ell}\frac{1}{\kappa}m_{r}\mathrm{d}\log z^{r}.$ (3.27)
Proposition 3.4.
The 1-form (3.27) is a left unity of the algebra structure defined by (3.26) on $T^{*}\mathcal{M}$.
Proof.
From the definition of the multiplication (3.26) it follows that
$\omega_{e}\cdot\mathrm{d}t^{\beta}=\frac{\kappa}{d_{\beta}}~{}\nabla_{g}^{\omega_{e}}\mathrm{d}t^{\beta}=\frac{\kappa}{d_{\beta}}\nabla^{g}_{g^{\sharp}(\omega_{e})}\mathrm{d}t^{\beta}.$ (3.28)
Here $\nabla_{g}$ (resp. $\nabla^{g}$) is the contravariant (resp. covariant dual) Levi-Civita connection, and $g^{\sharp}$ is the canonical isomorphism
$g^{\sharp}\colon T^{*}\mathcal{M}\to T^{**}\mathcal{M}=T\mathcal{M},\quad\omega\mapsto g(\omega,\cdot).$ (3.29)
So it suffices to show that
$\frac{\kappa}{d_{\beta}}\nabla^{g}_{g^{\sharp}(\omega_{e})}\mathrm{d}t^{\beta}=\mathrm{d}t^{\beta}.$ (3.30)
Note that the null-transformation $\varphi_{0}$ transforms the metric $a$ defined in (3.2) to $g$, hence $x^{1},\dots,x^{\ell}$ are the flat coordinates of $g$. Since the relation between the coordinates $z^{1},\dots,z^{\ell}$ and $t^{1},\dots,t^{\ell}$ is independent of $\lambda$, in the following proof of the validity of (3.30) we can assume that $z^{i}$ and $t^{\alpha}$ are related with $x^{1},\dots,x^{\ell}$ via the null-transformation $\varphi_{0}$. From (2.43) we obtain
$\varphi_{0}^{*}\omega_{e}=-\frac{2}{\kappa}\pi i\mathrm{d}(\omega,\hat{\mathbf{x}}),$ (3.31)
it follows that
$\frac{\kappa}{d_{\beta}}\nabla^{g}_{g^{\sharp}(\omega_{e})}\mathrm{d}t^{\beta}\bigr{|}_{\lambda=0}$ $=-2\pi i\frac{1}{d_{\beta}}\sum_{s=1}^{\ell}m_{s}\nabla^{g}_{g^{\sharp}\mathrm{d}x^{s}}\left(\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}\mathrm{d}x^{p}\right)$
$=-2\pi i\frac{1}{d_{\beta}}\sum_{r,s=1}^{\ell}m_{s}\frac{1}{4\pi^{2}}(\mathrm{d}x^{r},\mathrm{d}x^{s})\nabla^{a}_{\frac{\partial}{\partial x^{r}}}\left(\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}\right)\mathrm{d}x^{p}$
$=\frac{1}{d_{\beta}}\left(\sum_{r=1}^{\ell}\frac{1}{2\pi i}\theta_{r}\frac{\partial}{\partial x^{r}}\left(\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}\right)\right)\mathrm{d}x^{p}.$
Here we used the relation (3.3) for the third equality. Note that $t^{\beta}\bigr{|}_{\lambda=0}$ is a Fourier polynomial of degree $d_{\beta}$ which does not depend on $\lambda$, so $\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}$ and
$\sum_{r=1}^{\ell}\frac{1}{2\pi i}~{}\theta_{r}\frac{\partial}{\partial x^{r}}\left(\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}\right)=\left.d_{\beta}\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}.$ (3.32)
Thus we arrive at
$\frac{\kappa}{d_{\beta}}\nabla^{g}_{g^{\sharp}(\omega_{e})}\mathrm{d}t^{\beta}\bigr{|}_{\lambda=0}=\left.\frac{\partial t^{\beta}}{\partial x^{p}}\right|_{\lambda=0}\mathrm{d}x^{p}=\mathrm{d}t^{\beta}\bigr{|}_{\lambda=0}.$ (3.33)
The proposition is proved. ∎
Corollary 3.5.
Denote $\omega_{e}=e_{\alpha}\mathrm{d}t^{\alpha}$, then the following relations hold true:
$e_{\alpha}g^{\alpha\beta}=\frac{1}{\kappa}d_{\beta}t^{\beta},\quad e_{\alpha}\Gamma_{\gamma}^{\alpha\beta}=\frac{1}{\kappa}d_{\beta}\delta_{\gamma}^{\beta}.$ (3.34)
Proof.
The second relation of (3.34) follows immediately from Proposition 3.4. To prove the first relation of (3.34) we can assume, as in the proof of Proposition 3.4, that the coordinates $t^{1},\dots,t^{\ell}$ and $x^{1},\dots,x^{\ell}$ are related via the null-transformation $\varphi_{0}$. Thus we have
$e_{\alpha}g^{\alpha\beta}$ $=g(\omega_{e},\mathrm{d}t^{\beta})=-\frac{2}{\kappa}\pi ig(\mathrm{d}(\omega,\hat{\mathbf{x}}),\mathrm{d}t^{\beta})$
$=\frac{\kappa}{2\pi i}\sum_{s=1}^{\ell}m_{s}(\mathrm{d}x^{r},\mathrm{d}x^{s})\frac{\partial t^{\beta}}{\partial x^{r}}=\frac{1}{\kappa}\sum_{s=1}^{\ell}\frac{\theta_{s}}{2\pi i}\frac{\partial t^{\beta}}{\partial x^{s}}=d_{\beta}t^{\beta}.$
The corollary is proved. ∎
Proposition 3.6 ([9], Appendix D).
The Christoffel symbols $\Gamma_{\gamma}^{\alpha\beta}$ are uniquely determined by the metric $g$ via the relations
$\frac{\partial g^{\alpha\beta}}{\partial t^{\gamma}}=\Gamma_{\gamma}^{\alpha\beta}+\Gamma_{\gamma}^{\beta\alpha},\quad g^{\alpha\epsilon}\Gamma_{\epsilon}^{\beta\gamma}=g^{\beta\epsilon}\Gamma_{\epsilon}^{\alpha\gamma}.$ (3.35)
Moreover, we have the following properties for the flat pencil of metrics:
$\eta^{\alpha\epsilon}\Gamma_{\epsilon}^{\beta\gamma}=\eta^{\beta\epsilon}\Gamma_{\epsilon}^{\alpha\gamma},\quad\frac{\partial\Gamma_{\gamma}^{\alpha\beta}}{\partial t^{\delta}}=\frac{\partial\Gamma_{\delta}^{\alpha\beta}}{\partial t^{\gamma}},\quad\Gamma_{\epsilon}^{\alpha\beta}\Gamma^{\epsilon\gamma}_{\delta}=\Gamma_{\epsilon}^{\alpha\gamma}\Gamma^{\epsilon\beta}_{\delta}.$ (3.36)
It follows from the first two relations of (3.36) that there exist functions $f^{\beta}$ such that
$\eta^{\alpha\epsilon}\frac{\partial^{2}f^{\beta}}{\partial t^{\epsilon}\partial t^{\gamma}}=\Gamma^{\alpha\beta}_{\gamma},$ (3.37)
where $\deg f^{\beta}=\kappa+d_{\beta}$
Lemma 3.7.
The Christoffel symbols satisfy the following relations:
$\Gamma_{\alpha}^{\beta\gamma}=\frac{d_{\gamma}}{d_{\beta}+d_{\gamma}}\frac{\partial g^{\beta\gamma}}{\partial t^{\alpha}}.$ (3.38)
Proof.
By multiplying $e_{\alpha}$ on both sides of the second relation of (3.35) and by using (3.34) we obtain
$\sum_{\epsilon}d_{\epsilon}t^{\epsilon}\Gamma_{\epsilon}^{\beta\gamma}=d_{\gamma}g^{\beta\gamma},$ (3.39)
so it follows from (3.37) that
$\sum_{\epsilon}d_{\epsilon}t^{\epsilon}\frac{\partial}{\partial t^{\epsilon}}\left(\eta^{\beta\nu}\frac{\partial f^{\gamma}}{\partial t^{\nu}}\right)=(d_{\gamma}+d_{\beta})\eta^{\beta\nu}\frac{\partial f^{\gamma}}{\partial t^{\nu}}=d_{\gamma}g^{\beta\gamma}.$ (3.40)
Thus we have
$(d_{\gamma}+d_{\beta})\eta^{\beta\nu}\frac{\partial^{2}f^{\gamma}}{\partial t^{\nu}\partial t^{\alpha}}=d_{\gamma}\frac{\partial g^{\beta\gamma}}{\partial t^{\alpha}},$ (3.41)
which leads to the validity of (3.38). The lemma is proved. ∎
Theorem 3.8.
[…]
Proof.
We need to show that the algebra is commutative and associative, and the operation of multiplication is compatible with the bilinear form. From (3.38) it follows that
$\frac{1}{d_{\gamma}}\Gamma^{\beta\gamma}_{\alpha}=\frac{1}{d_{\beta}}\Gamma^{\gamma\beta}_{\alpha},$ (3.42)
which leads to the commutativity of the algebra. Now by using the third relation of (3.36) we have
$\left(\operatorname{d\!}t^{\alpha}\cdot\operatorname{d\!}t^{\beta}\right)\cdot\operatorname{d\!}t^{\gamma}=\frac{\kappa}{d_{\beta}}\Gamma^{\alpha\beta}_{\xi}\operatorname{d\!}t^{\xi}\cdot\operatorname{d\!}t^{\gamma}=\frac{\kappa}{d_{\beta}}\frac{\kappa}{d_{\gamma}}\Gamma^{\alpha\beta}_{\xi}\Gamma^{\xi\gamma}_{\zeta}\operatorname{d\!}t^{\zeta}$
$\quad=\frac{\kappa}{d_{\alpha}}\frac{\kappa}{d_{\gamma}}\Gamma^{\beta\alpha}_{\xi}\Gamma^{\xi\gamma}_{\zeta}\operatorname{d\!}t^{\zeta}=\frac{\kappa}{d_{\alpha}}\frac{\kappa}{d_{\gamma}}\Gamma^{\beta\gamma}_{\xi}\Gamma^{\xi\alpha}_{\zeta}\operatorname{d\!}t^{\zeta}$
$\quad=\left(\operatorname{d\!}t^{\beta}\cdot\operatorname{d\!}t^{\gamma}\right)\cdot\operatorname{d\!}t^{\alpha}=\operatorname{d\!}t^{\alpha}\cdot\left(\operatorname{d\!}t^{\beta}\cdot\operatorname{d\!}t^{\gamma}\right),$
so the associativity of the algebra holds true. Finally, from the first relation of (3.36) we obtain
$\langle\operatorname{d\!}t^{\alpha}\cdot\operatorname{d\!}t^{\beta},\operatorname{d\!}t^{\gamma}\rangle=\frac{\kappa}{d_{\beta}}\eta^{\gamma\xi}\Gamma^{\alpha\beta}_{\xi}=\frac{\kappa}{d_{\beta}}\eta^{\alpha\xi}\Gamma^{\gamma\beta}_{\xi}=\frac{\kappa}{d_{\gamma}}\eta^{\alpha\xi}\Gamma^{\beta\gamma}_{\xi}=\langle\operatorname{d\!}t^{\alpha},\operatorname{d\!}t^{\beta}\cdot\operatorname{d\!}t^{\gamma}\rangle,$
which verifies the compatibility of the multiplication with the bilinear form. The proposition is proved. ∎
The Frobenius algebra on $T^{*}\mathcal{M}$ induces a Frobenius algebra structure on $T(M\setminus D)$ via the isomorphism
$\eta^{\sharp}\colon T^{*}\mathcal{M}\to T^{**}\mathcal{M}=T\mathcal{M},\quad\omega\mapsto\eta(\omega,\cdot),$ (3.43)
where $D$ is define by (3.13). Explicitly, the Frobenius algebra structure $(\cdot,\langle\cdot\,,\cdot\rangle,e)$ on $T(M\setminus D)$ is defined by
$X\cdot Y=\eta^{\sharp}\left(\eta^{\flat}(X)\cdot\eta^{\flat}(Y)\right),\quad X,Y\in\rm{Vect}(\mathcal{M}),$ (3.44)
with the bilinear form
$\langle X,Y\rangle=\eta^{-1}(X,Y)=\eta(\eta^{\flat}(X),\eta^{\flat}(Y))$ (3.45)
and the unity
$e=\eta^{\sharp}(\omega_{e}),$ (3.46)
here $\eta^{\flat}$ is the inverse map of $\eta^{\sharp}$. From (3.14), (3.20) and (3.27) it follows that
$\mathcal{L}_{E}e=-\kappa e.$
3.4. The generalized Frobenius manifold structure on $\mathcal{M}\setminus D$
The notion of a generalized Frobenius manifold is a certain generalization of Dubrovin’s definition of Frobenius manifolds [9], it satisfies all axioms of the definition of a Frobenius manifold but the flatness of the unit vector field [22].
Definition 3.2 ([9, 22]).
[…]
  1. FM1.
[…]
  2. FM2.
    Define a $3$-tensor $c$ by $c(u,v,w)=\langle u\cdot v,w\rangle$, then the $4$-tensor
    $\nabla c:(W,X,Y,Z)\to(\nabla_{W}c)(X,Y,Z)$ (3.47)
    is symmetric w.r.t. $W,X,Y,Z\in\rm{Vect}(M)$.
  3. FM3.
    There exist an Euler vector field $E$ on $M$ such that $\nabla\nabla E=0$, $\nabla E$ is diagonalizable and
    $\mathcal{L}_{E}(X\cdot Y)$ $=(\mathcal{L}_{E}X)\cdot Y+X\cdot(\mathcal{L}_{E}Y)+X\cdot Y.$ (3.48)
    $\mathcal{L}_{E}\langle X,Y\rangle$ $=\langle\mathcal{L}_{E}X,Y\rangle+\langle X,\mathcal{L}_{E}Y\rangle+(2-d)\langle X,Y\rangle.$ (3.49)
    Here $\nabla$ is the Levi-Civita connection of the flat metric, and $d$ is a constant whic is called the charge of $M$.
Let us show that the Frobenius algebra on $T(\mathcal{M}\setminus D)$ that is defined in the last subsection yields a generalized Frobenius manifold structure on $\mathcal{M}\setminus D$ with the Euler vector field
$\tilde{E}=\frac{1}{\kappa}E=\sum_{\alpha=1}^{\ell}\frac{d_{\alpha}}{\kappa}\frac{\partial}{\partial t^{\alpha}}$ (3.50)
and with charge $d=1$. The validity of the axiom FM1 follows from the flatness of the metric $\eta$, and that of the axiom FM2 follows from the relation (3.42) and the second relation of (3.36). To check the quasi-homogeneity of the Frobenius manifold structure required by the axiom FM3, we need to verify equations (3.48) and (3.49) which can be represented in the flat coordinates $t^{1},\dots,t^{\ell}$ as follows:
$\partial_{\tilde{E}}c_{\alpha\beta}^{\gamma}$ $=\left(1-\frac{d_{\alpha}}{\kappa}-\frac{d_{\beta}}{\kappa}+\frac{d_{\gamma}}{\kappa}\right)c_{\alpha\beta}^{\gamma},$ (3.51)
00 0 $=\left(1-\frac{d_{\alpha}}{\kappa}-\frac{d_{\beta}}{\kappa}\right)\eta_{\alpha\beta}.$ (3.52)
Here
$c_{\alpha\beta}^{\gamma}=\frac{\kappa}{d_{\rho}}\eta_{\alpha\nu}\eta_{\beta\rho}\eta^{\gamma\zeta}\Gamma^{\nu\rho}_{\zeta}$ (3.53)
are the structure constants of the Frobenius algebra
$\frac{\partial}{\partial t^{\alpha}}\cdot\frac{\partial}{\partial t^{\beta}}=c_{\alpha\beta}^{\gamma}\frac{\partial}{\partial t^{\gamma}},$ (3.54)
and $(\eta_{\alpha\beta})=(\eta^{\alpha\beta})^{-1}$ with $\eta^{\alpha\beta}=\eta(\operatorname{d\!}t^{\alpha},\operatorname{d\!}t^{\beta})$. The validity of the identities follows from (3.20) and the fact that $\deg\Gamma^{\alpha\beta}_{\gamma}=d_{\alpha}+d_{\beta}-d_{\gamma}$. Thus we have the following theorem.
Theorem 3.9 (Main).
Suppose $\{z^{1},\dots,z^{\ell}\}$ be a set of pencil generators associated with an irreducible reduced root system $R$ and a fixed weight $\omega$, then there exists a generalized Frobenius manifold structure of charge $d=1$ on $\mathcal{M}\setminus D$, of which the flat metric is given by $\eta$ and the multiplication is defined by (3.44). Moreover, the unit vector field $e$ and the Euler vector field $\tilde{E}$ of this generalized Frobenius are given by (3.46) and (3.50) respectively, and the intersection form coincides with $g$.
Proof.
We only need to show that the intersection form of the generalized Frobenius manifold $\mathcal{M}\setminus D$ coincides with $g$, i.e.,
$g(\operatorname{d\!}t^{\alpha},\operatorname{d\!}t^{\beta})=i_{\tilde{E}}(\operatorname{d\!}t^{\alpha}\cdot\operatorname{d\!}t^{\beta}).$ (3.55)
Indeed, by using Lemma 3.7 we obtain
$i_{\tilde{E}}(\mathrm{d}t^{\alpha}\cdot\mathrm{d}t^{\beta})=\frac{\kappa\tilde{E}^{\gamma}}{d_{\beta}}\Gamma^{\alpha\beta}_{\gamma}=\sum_{\gamma}\frac{d_{\gamma}t^{\gamma}}{d_{\beta}}\Gamma^{\alpha\beta}_{\gamma}=\sum_{\gamma}\frac{d_{\gamma}t^{\gamma}}{d_{\alpha}+d_{\beta}}\frac{\partial g^{\alpha\beta}}{\partial t^{\gamma}}=g^{\alpha\beta}.$ (3.56)
Here we used the fact that $\deg g^{\alpha\beta}=d_{\alpha}+d_{\beta}$. The theorem is proved. ∎
Note that the flat coordinates $t^{1},\dots,t^{\ell}$ as functions of $z^{1},\dots,z^{\ell}$ live on the universal covering $\widehat{\mathcal{M}\setminus D}$ of $\mathcal{M}\setminus D$, hence $\pi_{1}(\mathcal{M}\setminus D)$ acts on those functions. Denote the stable subgroup of this action by
$H=\{\gamma\in\pi_{1}(\mathcal{M}\setminus D)\mid\gamma^{*}t^{\alpha}(p)=t^{\alpha}(p),~{}\forall p\in\widehat{\mathcal{M}\setminus D}\},$ (3.57)
then $H$ corresponds to a covering space $\overline{\mathcal{M}}$ of $\mathcal{M}\setminus D$. Concretely, we may define an equivalence relation $\sim$ on $\widehat{\mathcal{M}\setminus D}$ by
$p\sim p^{\prime}\iff\pi(p)=\pi(p^{\prime}),\quad t^{\alpha}(p)=t^{\alpha}(p^{\prime}),\quad\alpha=1,\dots,\ell.$ (3.58)
Here $\pi\colon\widehat{\mathcal{M}\setminus D}\to\mathcal{M}\setminus D$ is the canonical projection. Then we have
¯$\overline{\mathcal{M}}=\widehat{\mathcal{M}\setminus D}/\sim.$ (3.59)
We are to show in Corollary 4.3 that $t^{1},\dots,t^{\ell}$ are coordinates of $\overline{\mathcal{M}}$. Therefore, the generalized Frobenius manifold structure that we constructed above is defined on $\overline{\mathcal{M}}$ instead of on $\mathcal{M}\setminus D$. If there is no ambiguity, we usually identify $\overline{\mathcal{M}}$ with $\mathcal{M}\setminus D$ since they are locally isomorphic. However, when we discuss the monodromy group of the generalized Frobenius manifold structure in the next section, it is necessary to view this generalized Frobenius manifold as $\overline{\mathcal{M}}$. We will prove in the next section that $z^{1},\dots,z^{\ell}$ are polynomials of $t^{1},\dots,t^{\ell}$, so we have the canonical projection
$\overline{\pi}:~{}\overline{\mathcal{M}}\to\mathcal{M}\setminus D.$
Since $\overline{\mathcal{M}}$ is constructed from the root system $R$, the weight $\omega$ and the pencil generators $z^{1},\dots,z^{\ell}$, we will also denote this generalized Frobenius manifold by
¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega,\{z^{1},\dots,z^{\ell}\}).$ (3.60)
If there is no confusion on the selection of the pencil generators, we simply write it as
¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega).$ (3.61)
4. Polynomial Property and the Monodromy Groups
In this section, we consider the monodromy group of the generalized Frobenius manifold ¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega)$. Note that the monodromy groups of the Frobenius manifold structures that are constructed in [9, 12] on the orbit spaces of the Coxeter groups and the extended affine Weyl groups coincide with the underlined groups. In the present case the monodromy group $\mathrm{Mono}(\overline{\mathcal{M}})$ can be embedded into the affine Weyl group $W_{a}(R)$. More precisely, we will show that the image of this embedding satisfies
$\mathrm{Mono}(\overline{\mathcal{M}})~{}\leq~{}\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell},$ (4.1)
where
$\mathrm{Stab}(\omega)=\{\sigma\in W(R)\mid\sigma(\omega)=\omega\}.$ (4.2)
To this end, we first need to prove Theorem 4.2 which states that for the set $\{z^{1},\dots,z^{\ell}\}$ of pencil generators the functions $z^{i}$ are polynomials of the flat coordinates $t^{1},\dots,t^{\ell}$.
4.1. Periods of the generalized Frobenius manifold
Let us consider flat coordinates of the flat metric $g+\lambda\eta$ of $\overline{\mathcal{M}}$. They satisfy the equations
$\nabla_{g+\lambda\eta}\mathrm{d}u(\lambda)=0$ (4.3)
and are called periods of the generalized Frobenius manifold [9, 22]. Denote
$M_{\lambda}=(\overline{\mathcal{M}}\times\mathbb{C}^{\times})\setminus{\Sigma_{\lambda}},$ (4.4)
where
${\Sigma}_{\lambda}=\{(z,\lambda)\in M\times\mathbb{C}^{\times}\mid\det(g+\lambda\eta)(z)=0\}.$
Then solutions of the equation (4.3) live on
$\mathscr{Q}(\widehat{M_{\lambda}})=\mathscr{O}(\widehat{M_{\lambda}})/\mathbb{C}.$
In the coordinates $(z^{1},\dots,z^{\ell})$, the equation (4.3) reads
$\frac{\partial^{2}u}{\partial z^{i}\partial z^{j}}-\Gamma_{\lambda,ij}^{k}\frac{\partial u}{\partial z^{k}}=0,$ (4.5)
Here $\Gamma_{\lambda,ij}^{k}$ are the Christoffel symbols of the Levi-Civita connection of $g+\lambda\eta$. Note that the $\varphi_{\lambda}$-transformation (3.1) transforms the constant metric $a$ defined in (3.2) to $g+\lambda\eta$, hence $x^{1},\dots,x^{\ell}$ are flat coordinates of $g+\lambda\eta$. By the definition of proper generators, there exist Laurant polynomials $Z_{1},\dots,Z_{\ell}$ such that
$z^{r}=\mathrm{e}^{-2\pi i\theta_{r}c}Z_{r}\left(\mathrm{e}^{2\pi i(x^{1}+\theta_{1}c)},\dots,\mathrm{e}^{2\pi i(x^{\ell}+\theta_{\ell}c)}\right).$ (4.6)
Therefore, the refined flat coordinates
$\check{x}^{r}(z,\lambda):=x^{r}+\theta_{r}c=x^{r}+\frac{\theta_{r}}{\kappa}\log\lambda,\quad r=1,\dots,\ell$ (4.7)
depend only on $z^{j}\mathrm{e}^{2\pi i\theta_{j}c},\,j=1,\dots,\ell$. As functions of $z^{1},\dots,z^{\ell}$ and $\lambda$, they are homogeneous of degree zero, i.e., they satisfy the equation
$\kappa\lambda\frac{\partial u}{\partial\lambda}+\sum_{j=1}^{\ell}\theta_{j}z^{j}\frac{\partial u}{\partial z^{j}}=0.$ (4.8)
We denote by $\mathcal{S}_{\lambda}$ the space of periods of $\overline{\mathcal{M}}$ which are homogeneous of degree zero, then $\dim\mathcal{S}_{\lambda}=\ell$, and
$\mathcal{S}_{\lambda}=\mathrm{span}_{\mathbb{C}}\left\{\check{x}^{1}(z,\lambda),\dots,\check{x}^{\ell}(z,\lambda)\right\}.$ (4.9)
Now we proceed to find another basis of $\mathcal{S}_{\lambda}$ that is associated with the flat coordinates $t^{1}(z),\dots,t^{\ell}(z)$ of the metric $\eta$. We denote $\rho=\frac{1}{\lambda}$.
Proposition 4.1.
For each homogeneous flat coordinate $t^{\alpha}=t^{\alpha}(z)$ of the metric $\eta$ with degree $d_{\alpha}$, there is a solution
$u=v^{\alpha}(z,\rho)=\rho^{d_{\alpha}/\kappa}h^{\alpha}(z,\rho)$ (4.10)
of the equations (4.5) and (4.8), where $h^{\alpha}(z,\rho)$ can be analytically continued to $\rho=0$ such that $h^{\alpha}(z,0)=t^{\alpha}(z)$.
Proof.
Recall that we made the assumption in Remark 3.1 that $d_{\alpha}>0$, so from (3.20) we have $0<d_{\alpha}<\kappa$. Let us looking for a solution of (4.5) and (4.8) of the form $u(z,\rho)=\rho^{d_{\alpha}/\kappa}h(z,\rho)$ satisfying the initial condition $h(z,0)=t^{\alpha}(z)$. Denote $\xi_{j}=\frac{\partial h}{\partial z^{j}}$, then these equations take the form
$\frac{\partial\xi_{j}}{\partial z^{i}}-\Gamma_{ij}^{k}(\rho)\xi_{k}$ $=0,$ (4.11)
$-\kappa\rho\frac{\partial\xi_{j}}{\partial\rho}+\sum_{r=1}^{\ell}\theta_{r}z^{r}\frac{\partial\xi_{j}}{\partial z^{r}}$ $=(d_{\alpha}-\theta_{j})\xi_{j},$ (4.12)
where $\Gamma_{ij}^{k}(\rho)=\Gamma_{1/\rho,ij}^{k}$, which are the Christoffel symbols of the Levi-Civita connection of $\eta+\rho g$. By using the equations (4.11) we can rewrite the equations (4.12) as
$\kappa\rho\frac{\partial\vec{\xi}}{\partial\rho}=\left(A(\rho)-d_{\alpha}\right)\vec{\xi},$ (4.13)
where
$\vec{\xi}=\vec{\xi}(\rho)=(\xi_{1},\cdots,\xi_{\ell})^{T},\quad A_{j}^{k}(\rho)=\sum_{r=1}^{\ell}\theta_{r}z^{r}\Gamma_{rj}^{k}(\rho)+\theta_{j}\delta_{j}^{k}.$
Expand $A(\rho)$ and $\zeta{\xi}(\rho)$ into power series of $\rho$ as follows:
$A(\rho)=A+\rho A_{1}+\rho^{2}A_{2}+\dots,\quad\vec{\xi}(\rho)=\vec{\xi}_{0}+\rho\vec{\xi}_{1}+\rho^{2}\vec{\xi}_{2}+\dots,$
here $A=A(0)=(A^{k}_{j}(0))$ coincides with the matrix $A$ defined in (3.23), and
$\vec{\xi}_{0}=\left(\frac{\partial t^{\alpha}}{\partial z^{1}},\dots,\frac{\partial t^{\alpha}}{\partial z^{\ell}}\right)^{T},$ (4.14)
then the equation (4.13) is equivalent to
$(A-d_{\alpha})\vec{\xi}_{0}=\mathbf{0},\quad(A-d_{\alpha}-\kappa)\vec{\xi}_{1}=-A_{1}\vec{\xi}_{0},$ (4.15)
$(A-d_{\alpha}-2\kappa)\vec{\xi}_{2}=-A_{2}\vec{\xi}_{0}-A_{1}\vec{\xi}_{1},\dots$ (4.16)
From (3.24) we know that the first equation of (4.15) is satisfied by $\vec{\xi}_{0}$ that is defined by (4.14). Since the eigenvalues of $A$ are positive and less than $\kappa$, we can solve the other linear equations of (4.15) and (4.16) to obtain a unique solution $\vec{\xi}_{1},\vec{\xi}_{2},\dots$ recursively. From the fact that $\vec{\xi}_{0}$ gives a solution of the equations (4.11), and from the compatibility of of the equations of (4.11) and (4.12) we know that $\vec{\xi}_{1},\vec{\xi}_{2},\dots$ yield a solution of the equations (4.11) and (4.12), thus we proved the proposition. ∎
4.2. The polynomiality property
The periods $v^{1}(z,\rho),\cdots,v^{\ell}(z,\rho)$ of $\overline{\mathcal{M}}$ given by Proposition 4.1 are linearly independent, so they yield another basis of $\mathcal{S}_{\lambda}$. Thus there exist constants $c_{\alpha}^{\beta}$ and $s^{\beta}$ such that
$\check{x}^{\beta}(z,\rho)=c_{\alpha}^{\beta}v^{\alpha}(z,\rho)+s^{\beta}.$ (4.17)
Since the functions $Z_{r}=Z_{r}(\mathrm{e}^{2\pi i\check{x}^{1}},\dots,\mathrm{e}^{2\pi i\check{x}^{\ell}})$ given by (4.6) are holomorphic functions of $\check{x}^{1},\dots,\check{x}^{\ell}$, they are also holomorphic functions of $v^{1}(z,\rho),\dots,v^{\ell}(z,\rho)$. Hence we have
$z^{r}=\rho^{-\theta_{r}/\kappa}Z_{r}=\rho^{-\theta_{r}/\kappa}\sum_{n_{1},\dots,n_{\ell}\in\mathbb{Z}_{\geq 0}}a_{n_{1},\dots,n_{\ell}}(v^{1}(z,\rho))^{n_{1}}\cdots(v^{\ell}(z,\rho))^{n_{\ell}},$ (4.18)
where $a_{n_{1},\dots,n_{\ell}}$ are some constants.
Theorem 4.2.
For each $r=1,\cdots,\ell$, $z^{r}$ is a polynomial of $t^{1},\cdots,t^{\ell}$.
Proof.
It follows from Proposition 4.1 that
$v^{\alpha}(z,\rho)=\rho^{d_{\alpha}/\kappa}t^{\alpha}(z)+o(\rho^{d_{\alpha}/\kappa}),\quad\rho\to 0.$
Thus we have
$z^{r}=\rho^{-\theta_{r}/\kappa}\sum_{n_{1},\dots,n_{\ell}\in\mathbb{Z}_{\geq 0}}a_{n_{1},\dots,n_{\ell}}\left(\rho^{(n_{1}d_{1}+\dots+n_{\ell}d_{\ell})/\kappa}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}+o(\rho^{(n_{1}d_{1}+\dots+n_{\ell}d_{\ell})/\kappa})\right).$ (4.19)
Let us denote
$\theta=\inf\{n_{1}d_{1}+\dots+n_{\ell}d_{\ell}~{}\big{|}~{}a_{n_{1},\dots,n_{\ell}}\not=0\}.$ (4.20)
Suppose $\theta>\theta_{r}$, then it follows from (4.19) that
$z^{r}=\rho^{-\theta_{r}/\kappa}o(\rho^{\theta_{r}/\kappa})=o(1),\quad\rho\to 0,$ (4.21)
this is not true due to our definition of proper generators. On the other hand, if $\theta<\theta_{r}$, then from (4.19) we know that
$z^{r}=\rho^{(\theta-\theta_{r})/\kappa}\sum_{n_{1}d_{1}+\dots+n_{\ell}d_{\ell}=\theta}a_{n_{1},\dots,n_{\ell}}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}+o(\rho^{(\theta-\theta_{r})/\kappa}),\quad\rho\to 0.$ (4.22)
Since $(\theta-\theta_{r})/\kappa<0$, we must have
$\sum_{n_{1}d_{1}+\cdots+n_{\ell}d_{\ell}=\theta}a_{n_{1},\cdots,n_{\ell}}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}=0.$
However, since $t^{1},\cdots,t^{\ell}$ are functionally independent, all those coefficients $a_{n_{1},\cdots,n_{\ell}}$ must vanish. This contradicts with the definition (4.20) of $\theta$. Thus we have $\theta=\theta_{r}$, and it follows that
$z^{r}=\sum_{n_{1}d_{1}+\cdots+n_{\ell}d_{\ell}=\theta_{r}}a_{n_{1},\cdots,n_{\ell}}(t^{1})^{n_{1}}\cdots(t^{\ell})^{n_{\ell}}.$ (4.23)
In particular, $z^{r}$ is a polynomial of $t^{1},\cdots,t^{\ell}$. The theorem is proved. ∎
Remark 4.1.
This theorem shows that the proper generators are polynomials of the flat coordinates of $\eta$. It deserves to note that this property also holds true for the Frobenius manifold structures defined on the orbit spaces of Coxeter groups [9] and on the extended affine Weyl groups [12, 14]. For the Frobenius manifold structures constructed in [9] and [12], the flat coordinates $t^{\alpha}$ are also polynomials of the invariant polynomials $z^{j}$’s, while for the ones constructed in [14], $t^{\alpha}$ may involve radical expressions of $z^{j}$’s. For the generalized Frobenius manifold structures obtained by the construction of the present paper, the flat coordinates $t^{\alpha}$ are usually complicated functions of $z^{j}$’s, and even not solvable by radicals for some cases, such as $\mathcal{M}(C_{\ell},\omega_{\ell})$ for $\ell\geq 5$.
From Theorem 4.2, one can verify that each $g^{\alpha\beta}$ is a rational function of $t^{1},\cdots,t^{\ell}$. Now let us look at properties of the potential $F$ of the generalized Frobenius manifold $\overline{\mathcal{M}}$, which is a function of the flat coordinates $t^{1},\dots,t^{\ell}$ and is defined by the relations
$\eta^{\gamma\nu}\frac{\partial^{3}F}{\partial t^{\alpha}\partial t^{\beta}\partial t^{\nu}}=c^{\gamma}_{\alpha\beta},\quad\mathcal{L}_{\tilde{E}}F=2F+\textrm{quadratic functions of}\ t^{1},\dots,t^{\ell},$
up to the addition of quadratic functions of $t^{1},\dots,t^{\ell}$, here the structure constants $c^{\gamma}_{\alpha\beta}$ of the Frobenius algebra are defined in (3.53). From the relations (3.38) and (3.53) it follows that we can choose the potential $F$ such that
$\eta^{\alpha\nu}\eta^{\beta\xi}\frac{\partial^{2}F}{\partial t^{\nu}\partial t^{\xi}}=\frac{\kappa}{d_{\alpha}+d_{\beta}}g^{\alpha\beta}.$ (4.24)
Therefore, all the second derivatives of $F$ are rational functions of $t^{1},\cdots,t^{\ell}$.
Corollary 4.3.
Define the covering space $\overline{\mathcal{M}}$ of $\mathcal{M}\setminus D$ as (3.59). Then $(t^{1},\cdots,t^{\ell})$ serves as global coordinates of $\overline{\mathcal{M}}$.
Proof.
To prove that $(t^{1},\cdots,t^{\ell})$ are global coordinates, it suffices to show that
$(t^{1},\cdots,t^{\ell})\colon\overline{\mathcal{M}}\to\mathbb{C}^{\ell}$ (4.25)
is injective. Suppose $p,p^{\prime}\in\overline{\mathcal{M}}$ satisfy $t^{\alpha}(p)=t^{\alpha}(p^{\prime}),~{}\alpha=1,\cdots,\ell$, then by using Theorem 4.2 we know that
$z^{r}(t^{1}(p),\cdots,t^{\ell}(p))=z^{r}(t^{1}(p^{\prime}),\cdots,t^{\ell}(p^{\prime})).$
This implies that $\overline{\pi}(p)=\overline{\pi}(p^{\prime})$, where $\overline{\pi}:\overline{\mathcal{M}}\to\mathcal{M}\setminus D$ is the canonical projection. Therefore $p\sim p^{\prime}$, i.e., they are identified on $\overline{\mathcal{M}}$. Hence (4.25) is injective, and we proved the corollary. ∎
In conclusion, we have the following commutative diagram:
[…]
In other words, upon taking coordinates $(t^{1},\cdots,t^{\ell})$ on $\overline{\mathcal{M}}$ and $(z^{1},\cdots,z^{\ell})$ on $\mathcal{M}\setminus D$, the projection $\overline{\pi}$ coincides with the natural map $\mathrm{Spec}(\mathbb{C}[t^{1},\cdots,t^{\ell}])\to\mathrm{Spec}(\mathbb{C}[z^{1},\cdots,z^{\ell}])$.
4.3. The Monodromy group of $\mathcal{M}(R,\omega)$
The definition of the monodromy group of a Frobenius manifold was given by Dubrovin in [9], and it can be extended directly to generalized Frobenius manifolds. Let us first recall this defintion.
Let $M$ be an $\ell$-dimensional generalized Frobenius manifold, of which the flat metric is $\eta$ and the intersection form is $g$, and let $t^{1},\cdots,t^{\ell}$ be flat coordinates of $\eta$. Consider the system of equations
$\frac{\partial^{2}u}{\partial t^{\alpha}\partial t^{\beta}}=\Gamma_{\alpha\beta}^{\epsilon}\frac{\partial u}{\partial t^{\epsilon}},\quad\alpha,\beta=1,\dots,\ell$ (4.26)
for the flat coordinates of $g$. Here $\Gamma_{\alpha\beta}^{\epsilon}$ are the Christoffel symbols of the Levi-Civita connection of $g$. Denote $\Sigma=\{t\in M~{}|~{}\det g(t)=0\}$, then the solutions $u$ live on the universal covering $\widehat{M\setminus\Sigma}$ of $M\setminus\Sigma$. Define the solution space
[…]
then we know that $\dim\mathcal{S}=n$, and there exist flat coordinates $u^{1},\cdots,u^{\ell}$ of $g$ such that
$\mathcal{S}=\mathrm{span}\{u^{1},\cdots,u^{\ell}\}.$ (4.28)
For any loop $\gamma\in\pi_{1}(M\setminus\Sigma)$, we have a transformation
$\gamma^{*}u^{\alpha}=A_{\beta}^{\alpha}(\gamma)u^{\beta}+B^{\alpha}(\gamma),\quad\alpha=1,\dots,\ell$ (4.29)
on $\mathcal{S}$ by analytic continuation. Denote by $E^{\ell}$ the dual space of $\mathcal{S}$, then it has coordinates $u^{1},\cdots,u^{\ell}$. Hence the transformation (4.29) yields a transformation on $E^{\ell}$, which we also denote by $\gamma^{*}$. Since $E^{\ell}$ is an $\ell$-dimensional complexified Euclidean space $\mathbb{R}^{\ell}\otimes\mathbb{C}$ equipped with the metric $g$, and the action by $\pi_{1}(M\setminus\Sigma)$ preserves the metric $g$, this transformation on $E^{\ell}$ is an isometry. Hence we have obtained a representation
$r\colon\pi_{1}(M\setminus\Sigma)\to\mathrm{Iso}(E^{\ell})\cong O(\ell)\rtimes\mathbb{C}^{\ell},\quad\gamma\mapsto\gamma^{*}.$ (4.30)
Definition 4.1 ([9], Appendix G.).
The image of $r$ given in (4.30) is called the monodromy group of $M$, which is denoted by
$\mathrm{Mono}(M)=r(\pi_{1}(\mathcal{M}\setminus\Sigma))\leq O(\ell)\rtimes\mathbb{C}^{\ell}.$ (4.31)
Now we proceed to compute the monodromy group of the generalized Frobenius manifold ¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega)$. Let
$\overline{\Sigma}=\{t\in\overline{\mathcal{M}}\mid\det\left(g^{\alpha\beta}(t)\right)=0\},$
and $\mathcal{U}(\mathcal{M})$ be the universal covering space of ¯$\overline{\mathcal{M}}\setminus\overline{\Sigma}$. The flat coordiantes of $g$ are $x^{1}\bigr{|}_{\lambda=0},\cdots,x^{\ell}\bigr{|}_{\lambda=0}$, and they are functions on $\mathcal{U}(\mathcal{M})$. Moreover, $x^{1}\bigr{|}_{\lambda=0},\cdots,x^{\ell}\bigr{|}_{\lambda=0}$ are also coordinates of $V\otimes\mathbb{C}$, so we can view the complexified Euclidean space $E^{\ell}$ as $V\otimes\mathbb{C}$. Hence we obtain a representation
$r\colon\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma})\to\mathrm{Iso}(V\otimes\mathbb{C})\cong O(\ell)\rtimes\mathbb{C}^{\ell},\quad\gamma\mapsto\gamma^{*},$ (4.32)
where
$\gamma^{*}x^{i}\bigr{|}_{\lambda=0}=A_{j}^{i}(\gamma)x^{j}\bigr{|}_{\lambda=0}+B^{i}(\gamma).$ (4.33)
Here the image $r(\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma}))$ is the monodromy group of $\overline{\mathcal{M}}$. However, this group is difficult to compute, and we will instead consider a group bigger than it.
Recall that the covering map
$\overline{\pi}\colon\overline{\mathcal{M}}\to\mathcal{M}\setminus D$ (4.34)
is nondegerated, so from
$\det\left(g^{\alpha\beta}(t)\right)=\det\left(g^{ij}(z)\right)\det\left(\frac{\partial t^{\alpha}}{\partial z^{j}}\right)$ (4.35)
we know that $\left(g^{\alpha\beta}(t)\right)$ degenerates if and only if $\left(g^{ij}(z(t)\right)$ degenerates. In other words, denote
$\Sigma^{\prime}=\{z\in\mathcal{M}\setminus D\mid\det\left(g^{ij}(z)\right)=0\},$
then we have $(\overline{\pi})^{-1}(\Sigma^{\prime})=\overline{\Sigma}$, and the map
$\overline{\pi}:\overline{\mathcal{M}}\setminus\overline{\Sigma}\to\mathcal{M}\setminus(D\cup\Sigma^{\prime})$ (4.36)
is also a covering map, hence
$\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma})\leq\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})).$ (4.37)
Since ¯$\overline{\mathcal{M}}\setminus\overline{\Sigma}$ is a covering space of $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, they share the same universal covering space $\mathcal{U}(\mathcal{M})$, and both $\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma})$ and $\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$ act on $\mathcal{U}(\mathcal{M})$. Note that $g^{ij}(z)$ are polynomials of $z^{1},\cdots,z^{\ell}$, so in particular they are single valued functions on $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$. This implies that the action of $\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$ on $\mathcal{U}(\mathcal{M})$ also leaves the metric $g$ invariant, so the representation (4.32) can be extended to a representation
$r\colon\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))\to\mathrm{Iso}(V\otimes\mathbb{C})\cong O(\ell)\rtimes\mathbb{C}^{\ell}.$ (4.38)
Thus it follows from (4.37) that
$\mathrm{Mono}(\overline{\mathcal{M}})=r(\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma}))~{}\leq~{}r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))),$ (4.39)
i.e., the monodromy group of $\overline{\mathcal{M}}$ is a subgroup of $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$. We are to show in what follows that $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$ is isomorphic to $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$.
Recall that
$\varphi_{0}\colon V\otimes\mathbb{C}\to\mathcal{M}$ (4.40)
is a quotient map. By eliminating a divisor, we obtain a covering map
$\varphi_{0}\colon V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})\to\mathcal{M}\setminus(D\cup\Sigma^{\prime}).$ (4.41)
Since $\mathcal{U}(\mathcal{M})$ is the universal covering space of $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, it is also a covering space of $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$, and the covering map is given by
$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}\colon\mathcal{U}(\mathcal{M})\to V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime}),\quad p\mapsto(x^{1}\bigr{|}_{\lambda=0}(p),\cdots,x^{\ell}\bigr{|}_{\lambda=0}(p)).$ (4.42)
In conclusion, we have the following commutative diagram of covering spaces:
$\mathcal{U}(\mathcal{M})$¯$\overline{\mathcal{M}}\setminus\overline{\Sigma}$$\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$\mathrm{pr}_{1}$$\overline{\pi}$$\mathrm{pr}_{2}$$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$$\varphi_{0}$
All the maps in the diagram are covering maps. From the property of universal covering spaces, we have
$\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))=\mathrm{Deck}\left(\frac{\mathcal{U}(\mathcal{M})}{\mathcal{M}\setminus(D\cup\Sigma^{\prime})}\right).$ (4.43)
Here $\mathrm{Deck}(X/Y)$ is the deck transformation group on $X$ over $Y$, where $X$ is a covering space over $Y$. Relation (4.43) implies that, every element $\gamma\in\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$ induces a deck transformation on $\mathcal{U}(\mathcal{M})$. We have the following lemma for $r(\gamma)$.
Lemma 4.4.
Let $\gamma\in\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$, and let us view $\gamma$ as a deck transformation on $\mathcal{U}(\mathcal{M})$. Then $p=r(\gamma)$ is the unique deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$ such that
$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}\circ\gamma=p\circ\hat{\mathbf{x}}\bigr{|}_{\lambda=0}.$ (4.44)
In other words, we have the following commutative diagram.
$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.$\mathcal{U}(\mathcal{M})$$\mathcal{U}(\mathcal{M})$$\exists!~{}p=r(\gamma)$$\varphi_{0}$$\varphi_{0}$$\gamma$$\hat{\mathbf{x}}\Big{|}_{\lambda=0}$$\hat{\mathbf{x}}\Big{|}_{\lambda=0}$
Proof.
Since $\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$ is surjective, the map $p$ satisfying (4.44) is uniquely determined. Moreover, since $r(\gamma)=\gamma^{*}$ is defined as analytic continuation on $\mathcal{U}(\mathcal{M})$, $p=r(\gamma)$ satisfies (4.44). Thus in order to prove the lemma, it suffices to verify that $r(\gamma)$ is a deck transformation over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, i.e., $\varphi_{0}\circ r(\gamma)=\varphi_{0}$. As $\gamma$ is a deck transformation over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, we have
$\varphi_{0}\circ\hat{\mathbf{x}}\bigr{|}_{\lambda=0}=\varphi_{0}\circ\hat{\mathbf{x}}\Big{|}_{\lambda=0}\circ\gamma=\varphi_{0}\circ r(\gamma)\circ\hat{\mathbf{x}}\bigr{|}_{\lambda=0}.$ (4.45)
Since $\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$ is surjective, it follows that
$\varphi_{0}=\varphi_{0}\circ r(\gamma).$
Therefore, $r(\gamma)$ is a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$. The lemma is proved. ∎
Proposition 4.5.
The group $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$ can be represented as
$r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))=\mathrm{Iso}(V\otimes\mathbb{C})\cap\mathrm{Deck}\bigg{(}\frac{V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})}{\mathcal{M}\setminus(D\cup\Sigma^{\prime})}\bigg{)}.$ (4.46)
The right hand side consists of all the isometries on $V\otimes\mathbb{C}$ whose restrictions on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ are deck transformations over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.
Proof.
For any $\gamma\in\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime}))$, $r(\gamma)$ is of course in $\mathrm{Iso}(V\otimes\mathbb{C})$. Moreover, by using Lemma 4.4, we know that $r(\gamma)$ is a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.
On the other hand, suppose $p$ be an isometry of $V\otimes\mathbb{C}$ and also be a deck transformation on $V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$ over $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$. By the general lifting lemma of covering spaces, we know that the deck transformation $p$ can be lifted to a deck transformation $\sigma:\mathcal{U}(\mathcal{M})\to\mathcal{U}(\mathcal{M})$, since
$p_{*}\left(\hat{\mathbf{x}}\bigr{|}_{\lambda=0}\right)_{*}\pi_{1}(\mathcal{U}(\mathcal{M}))=\{e\}.$ (4.47)
Therefore, we have the following commutative diagram:
$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$V\otimes\mathbb{C}\setminus\varphi_{0}^{-1}(D\cup\Sigma^{\prime})$$\mathcal{M}\setminus(D\cup\Sigma^{\prime})$.$\mathcal{U}(\mathcal{M})$$\mathcal{U}(\mathcal{M})$$p$$\varphi_{0}$$\varphi_{0}$$\sigma$$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$$\hat{\mathbf{x}}\bigr{|}_{\lambda=0}$
Thus by using Lemma 4.4, we have $p=r(\sigma)$. The proposition is proved. ∎
Remark 4.2.
Note that $z^{1},\cdots,z^{\ell}$ are coordinates on $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, so intuitively speaking, $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$ consists of those isometric transformations on $(x^{1}\bigr{|}_{\lambda=0},\cdots,x^{\ell}\bigr{|}_{\lambda=0})$ that preserve $(z^{1},\cdots,z^{\ell})$. Similarly, $r(\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma}))$ consists of those isometric transformations on $(x^{1}\bigr{|}_{\lambda=0},\cdots,x^{\ell}\bigr{|}_{\lambda=0})$ that preserve $(t^{1},\cdots,t^{\ell})$. From Theorem 4.2 we know that, if $t^{1},\cdots,t^{\ell}$ are preserved under some transformations, so are $z^{1},\cdots,z^{\ell}$. This fact implies the relation (4.39).
We are ready to present the following theorem.
Theorem 4.6.
For the generalized Frobenius manifold ¯$\overline{\mathcal{M}}=\mathcal{M}(R,\omega)$, the group $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))$ is isomorphic to a subgroup of $W_{a}(R)\cong W(R)\rtimes\mathbb{Z}^{\ell}$. More precisely,
$r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))\cong\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}.$ (4.48)
Here $\mathrm{Stab}(\omega)$ is the stabilizer of $\omega$ in $W(R)$.
Proof.
Due to Proposition 4.5, we only need to determine all the isometric actions on $x^{i}\bigr{|}_{\lambda=0}$ that preserve $\mathcal{M}\setminus(D\cup\Sigma^{\prime})$, i.e., preserve $z^{1},\cdots,z^{\ell}$. From (2.31) and (2.41) it follows that
$z^{j}\bigr{|}_{\lambda=0}=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{*}(\omega_{j}),\hat{\mathbf{x}})}.$ (4.49)
Here
$H_{j}=\{\sigma\in W(R)\mid(\omega_{j}-\sigma^{*}(\omega_{j}),\omega)=0\}$
is a subgroup of $W(R)$. For any given element $\tilde{\phi}\in\mathrm{Iso}(E^{\ell})$, we decompose it into linear and the affine parts as follows:
$\tilde{\phi}(\mathbf{t})=\phi(\mathbf{t})+v^{1}\alpha_{1}^{\vee}+\dots+v^{\ell}\alpha_{\ell}^{\vee}.$ (4.50)
It acts on the the Fourier polynomials $z^{j}$ by
$\tilde{\phi}^{*}z^{j}\big{|}_{\lambda=0}$ $=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{*}(\omega_{j}),\tilde{\phi}(\hat{\mathbf{x}}))}$ (4.51)
$=\frac{1}{N_{j}}\sum_{\sigma\in H_{j}}\mathrm{e}^{2\pi i(\sigma^{*}(\omega_{j}),\phi(\hat{\mathbf{x}})+v^{1}\alpha_{1}^{\vee}+\dots+v^{\ell}\alpha_{\ell}^{\vee})},$ (4.52)
so it preserves $z^{j}\bigr{|}_{\lambda=0}$, $j=1,\cdots,\ell$ if and only if $\phi$ induces a permutation on $H_{j}$ by right action $\sigma\to\sigma\phi$ and
$(\sigma^{*}(\omega_{j}),v^{1}\alpha_{1}^{\vee}+\cdots+v^{\ell}\alpha_{\ell}^{\vee})$
is always an integer for each $\sigma\in H_{j}$, $j=1,\cdots,\ell$. These conditions are equivalent to
$\phi\in\bigcap_{j=1}^{\ell}H_{j},\quad v^{1},\cdots,v^{\ell}\in\mathbb{Z}.$ (4.53)
This implies that
$r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))=\bigcap_{j=1}^{\ell}H_{j}\rtimes\mathbb{Z}^{\ell}.$ (4.54)
Now let us compute $\bigcap_{j=1}^{\ell}H_{j}$. Note that $\sigma\in H_{j}$, $j=1,\cdots,\ell$ if and only if
$(\omega_{j},\omega)=(\omega_{j},\sigma(\omega)),\quad j=1,\cdots,\ell.$ (4.55)
Since $\omega_{j},j=1,\cdots,\ell$ span $V\otimes\mathbb{C}$, (4.55) is equivalent to $\sigma(\omega)=\omega$. Therefore,
$\bigcap_{j=1}^{\ell}H_{j}=\mathrm{Stab}(\omega).$ (4.56)
It follows from (4.54) and (4.56) that $r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))=\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$. The theorem is proved. ∎
By using (4.39), we have the following corollary.
Corollary 4.7.
The monodromy group of the generalized Frobenius manifold $\mathcal{M}(R,\omega)$ is isomorphic to a subgroup of $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$.
It comes out a natural question to ask: is $\mathrm{Mono}(\mathcal{M}(R,\omega))$ exactly isomorphic to $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$? Unfortunately, the answer to this question is negative, and we will see counterexamples in Sec. 5. However, in some special cases, the isomorphism holds true. One simple case is when ¯$\overline{\mathcal{M}}=\mathcal{M}\setminus D$. In this case we have ¯$\overline{\mathcal{M}}\setminus\overline{\Sigma}\cong(\mathcal{M}\setminus D)\setminus\Sigma^{\prime}$, hence
$\mathrm{Mono}(\overline{\mathcal{M}})=r(\pi_{1}(\overline{\mathcal{M}}\setminus\overline{\Sigma}))=r(\pi_{1}(\mathcal{M}\setminus(D\cup\Sigma^{\prime})))~{}\cong~{}\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}.$ (4.57)
We will also see examples which belong to this case in Sec. 5.
5. Examples
[…]
We start with the simplest example.
Example 5.1 ($(R,\omega)=(A_{1},\omega_{1})$).
Let $R$ be the root system of type $A_{1}$, $V$ be the hyperplane of $\mathbb{R}^{2}$ spanned by
$\alpha_{1}=e_{1}-e_{2}.$
We take $\omega$ to be the fundamental weight $\omega_{1}=\frac{1}{2}\alpha^{\vee}_{1}=\frac{1}{2}\alpha_{1}$. Let us introduce the coordinate $x_{1}$ on $V$ by
$\mathbf{x}=c\omega_{1}+x_{1}\alpha^{\vee}_{1},\quad\mathbf{x}\in V.$
In this case $\kappa=(\omega_{1},\alpha_{1})=1$, hence
$\lambda=\mathrm{e}^{-2\pi ic}.$
The Weyl group $W(R)$ acting on $V$ is generated by the reflections
$\sigma_{1}(\mathbf{x})=\mathbf{x}-(\alpha^{\vee}_{1},\mathbf{x})\alpha_{1},\quad\mathbf{x}\in V.$
Then we have
$\sigma_{1}(\mathbf{x})=c\omega_{1}+(-c-x_{1})\alpha^{\vee}_{1},\quad x_{1}\mapsto-c-x_{1}.$
The invariant Fourier polynomial has the form
$Y_{1}=\mathrm{e}^{2\pi i\left(x_{1}+\frac{1}{2}c\right)}+\mathrm{e}^{-2\pi i\left(x_{1}+\frac{1}{2}c\right)},$
and we take
$y_{1}=e^{-\pi ic}Y_{1}=\mathrm{e}^{2\pi ix_{1}}+\lambda\mathrm{e}^{-2\pi ix_{1}}.$
We have
$(\operatorname{d\!}x_{1},\operatorname{d\!}x_{1})=(\alpha^{\vee}_{1},\alpha^{\vee}_{1})^{-1}=\frac{1}{2},$
and
$g^{11}_{\lambda}=\frac{1}{4\pi^{2}}(\operatorname{d\!}y_{1},\operatorname{d\!}y_{1})=\frac{1}{4\pi^{2}}\frac{\partial y_{1}}{\partial x_{1}}(\operatorname{d\!}x_{1},\operatorname{d\!}x_{1})\frac{\partial y_{1}}{\partial x_{1}}=-\frac{1}{2}y_{1}^{2}+2\lambda.$
Therefore, here $\{y_{1}\}$ is a pencil generator, and we obtain a flat pencil
$\eta^{11}=2,\quad g^{11}=-\frac{1}{2}y_{1}^{2}.$
Now we consider the flat coordinate $t_{1}=y_{1}/\sqrt{2}$ of $\eta$, then in the new coordinate the flat pencil has the form
$\eta^{11}=1,\quad g^{11}=-\frac{1}{2}t_{1}^{2}.$
Thus we obtain a generalized Frobenius manifold $\mathcal{M}(A_{1},\omega_{1})$ with potential
$F=-\frac{1}{24}t_{1}^{4}.$
The Euler vector field and the unit vector field are given by
$E=\frac{1}{2}t_{1}\frac{\partial}{\partial t_{1}},\quad e=-\frac{1}{t_{1}}\frac{\partial}{\partial t_{1}}.$
Moreover, the flat coordinate of $g$ is given by
$x^{1}\big{|}_{\lambda=0}=\frac{1}{2\pi i}\log y_{1}=\frac{1}{2\pi i}(\log(\sqrt{2}t_{1})).$
It follows that only translations $x_{1}\to x_{1}+n$, $n\in\mathbb{Z}$ preserve $t_{1}$. Hence the monodromy group of the generalized Frobenius manifold is given by
$\mathrm{Mono}(\mathcal{M}(A_{1},\omega_{1}))=\mathbb{Z}.$
We considered the construction of generalized Frobenius manifolds for different types of affine Weyl groups with appropriately chosen weights. In what follows we will give some representative examples, which are chosen based on the following points:
  1. 1.
    In Example 5.2, the basic generators $y^{1},\cdots,y^{\ell}$ are pencil generators. The covering space ¯$\overline{\mathcal{M}}\to\mathcal{M}\setminus D$ is trivial, thus
    $\mathrm{Mono}(\mathcal{M}(R,\omega))~{}\cong~{}\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}.$
  2. 2.
    In Example 5.3, the basic generators $y^{1},\cdots,y^{\ell}$ are also pencil generators. However, unlike the previous example, the monodromy group
    $\mathrm{Mono}(\mathcal{M}(R,\omega))~{}\cong~{}\{e\}\rtimes\mathbb{Z}^{\ell}$
    is a proper subgroup of $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$.
  3. 3.
    In Example 5.4, the basic generators $y^{1},y^{2},y^{3}$ are not pencil generators, and we need to construct pencil generators by simple translations on $y^{1},y^{2}$ and $y^{3}$, i.e.,
    $z_{1}=y_{1}+a_{1}\lambda,\quad z_{2}=y_{2}+a_{2}\lambda,\quad z_{3}=y_{3}.$
    Here $a_{1},a_{2}$ are some constants, and we find three possible values for them.
  4. 4.
    In Example 5.5, $y^{1},y^{2}$ are not pencil generators. The pencil generators are constructed in a more complicated way as follows:
    $z_{1}=y_{1}-6\lambda,\quad z_{2}=y_{2}-3\lambda y_{1}+12\lambda^{2}.$
  5. 5.
    In Example 5.6, we select the weight $\omega=\omega_{1}+\omega_{3}$ instead of a fundamental weight. This complicates the construction of pencil generators and the generalized Frobenius manifold structure.
These examples involve most issues that we may come across in the computations. Something interesting issues will also happen when we consider other types of affine Weyl groups. More details will be presented in our subsequent publications.
Example 5.2 ($(R,\omega)=(A_{2},\omega_{2})$).
We take the root system $R$ of type $A_{2}$. Let $V$ be the hyperplane of $\mathbb{R}^{3}$ spanned by
$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3}.$
We have the coroots and fundamental weights
$\alpha^{\vee}_{1}=\alpha_{1},\quad\alpha^{\vee}_{2}=\alpha_{2},\quad\omega_{1}=\frac{2}{3}\alpha^{\vee}_{1}+\frac{1}{3}\alpha^{\vee}_{2},\quad\omega_{2}=\frac{1}{3}\alpha^{\vee}_{1}+\frac{2}{3}\alpha^{\vee}_{2}.$
We choose $\omega=\omega_{2}$. Let us introduce the coordinates $x_{1},x_{2}$ of $V$ by
$\mathbf{x}=c\omega_{2}+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2},\quad\mathbf{x}\in V.$
The Weyl group $W(R)$ acting on $V$ is generated by the reflections
$\sigma_{j}(\mathbf{x})=\mathbf{x}-(\alpha^{\vee}_{j},\mathbf{x})\alpha_{j},\quad j=1,2,\ \mathbf{x}\in V.$
The invariant $\lambda$-Fourier polynomials are given by
$Y_{1}$ $=\mathrm{e}^{2\pi i\left(x_{1}+\frac{1}{3}c\right)}+\mathrm{e}^{2\pi i\left(x_{2}-x_{1}+\frac{1}{3}c\right)}+\mathrm{e}^{-2\pi i\left(x_{2}+\frac{2}{3}c\right)},$
$Y_{2}$ $=\mathrm{e}^{2\pi i\left(x_{2}+\frac{2}{3}c\right)}+\mathrm{e}^{-2\pi i\left(x_{1}+\frac{1}{3}c\right)}+\mathrm{e}^{2\pi i\left(x_{1}-x_{2}-\frac{1}{3}c\right)}.$
Since $\kappa=1$, we have
$\lambda=\mathrm{e}^{-2\pi ic}.$ (5.1)
We take
$y_{1}$ $=\lambda^{\frac{1}{3}}Y_{1}=\mathrm{e}^{2\pi ix_{1}}+\mathrm{e}^{2\pi i(x_{2}-x_{1})}+\lambda\mathrm{e}^{-2\pi ix_{2}},$
$y_{2}$ $=\lambda^{\frac{2}{3}}Y_{2}=\mathrm{e}^{2\pi ix_{2}}+\lambda\mathrm{e}^{-2\pi ix_{1}}+\lambda\mathrm{e}^{2\pi i(x_{1}-x_{2})}.$
The metric on $V$ that is induced from $\mathbb{R}^{3}$ is given by
$\left((\alpha^{\vee}_{i},\alpha^{\vee}_{j})\right)=\begin{pmatrix}2&-1\\ -1&2\end{pmatrix},\quad(\operatorname{d\!}x_{i},\operatorname{d\!}x_{j})={\begin{pmatrix}2&-1\\ -1&2\end{pmatrix}}^{-1}=\frac{1}{3}\begin{pmatrix}2&1\\ 1&2\end{pmatrix}.$
In the coordinates $y_{1},y_{2}$ we have
$\left(g_{\lambda}^{ij}\right)=\frac{1}{4\pi^{2}}\left((\operatorname{d\!}y_{i},\operatorname{d\!}y_{j})\right)=\frac{1}{4\pi^{2}}\sum_{k,r=1}^{2}\frac{\partial y_{i}}{\partial x_{k}}\frac{\partial y_{j}}{\partial x_{r}}(\operatorname{d\!}x_{k},\operatorname{d\!}x_{r})=\begin{pmatrix}-\frac{2}{3}y_{1}^{2}+2y_{2}&3\lambda-\frac{1}{3}y_{1}y_{2}\\[3.0pt] 3\lambda-\frac{1}{3}y_{1}y_{2}&2\lambda y_{1}-\frac{2}{3}y_{2}^{2}\end{pmatrix}.$
Then $y^{1},y^{2}$ are pencil generators, and the corresponding flat pencil is given by
$\left(\eta^{ij}\right)=\begin{pmatrix}0&3\\ 3&2y_{1}\end{pmatrix},\quad\left(g^{ij}\right)=\begin{pmatrix}-\frac{2}{3}y_{1}^{2}+2y_{2}&-\frac{1}{3}y_{1}y_{2}\\[3.0pt] -\frac{1}{3}y_{1}y_{2}&-\frac{2}{3}y_{2}^{2}\end{pmatrix}.$
The metric $\eta$ has the flat coordinates
$t_{1}=\frac{1}{3}y_{1},\quad t_{2}=y_{2}-\frac{1}{6}y_{1}^{2},\quad$
in which the metrics $\eta$ and $g$ have the form
$\left(\eta^{ij}\right)=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\quad\left(g^{ij}\right)=\begin{pmatrix}\frac{1}{9}\left(2t_{2}-3t_{1}^{2}\right)&-t_{1}t_{2}+\frac{1}{2}t_{1}^{3}\\[4.0pt] -t_{1}t_{2}+\frac{1}{2}t_{1}^{3}&-\frac{1}{6}\left(2t_{2}-3t_{1}^{2}\right)^{2}\end{pmatrix}.$
Thus we obtain a generalized Frobenius manifold $\mathcal{M}(A_{2},\omega_{2})$ with potential
$F=\frac{1}{18}t_{2}^{3}-\frac{1}{4}t_{1}^{2}t_{2}^{2}+\frac{1}{8}t_{1}^{4}t_{2}-\frac{3}{80}t_{1}^{6},$ (5.2)
the Euler vector field
$E=\frac{1}{3}t_{1}\frac{\partial}{\partial t_{1}}+\frac{2}{3}t_{2}\frac{\partial}{\partial t_{2}},$
and the unit vector field
$e=-\frac{2}{2t_{2}+3t_{1}^{2}}\frac{\partial}{\partial t_{1}}-\frac{6t_{1}}{2t_{2}+3t_{1}^{2}}\frac{\partial}{\partial t_{2}}.$
We also have $e^{\flat}=-\mathrm{d}\log y_{2}$. Note that in this example, not only $y_{1},y_{2}$ are polynomials of $t^{1}$ and $t^{2}$, but also $t^{1},t^{2}$ are polynomials of $y_{1}$ and $y_{2}$. This implies that the monodromy between $t^{1}$, $t^{2}$ and $y_{1}$, $y_{2}$ are trivial, i.e., ¯$\overline{\mathcal{M}}=\mathcal{M}\setminus D$. Therefore,
$\mathrm{Mono}(\mathcal{M}(A_{2},\omega_{2}))=\mathrm{Stab}(\omega_{2})\rtimes\mathbb{Z}^{2}=\langle\sigma_{1}\rangle\rtimes\mathbb{Z}^{2}\cong\mathbb{Z}_{2}\rtimes\mathbb{Z}^{2}.$
More explicitly, $\mathrm{Mono}(\mathcal{M}(A_{2},\omega_{2}))$ is generated by
$\sigma_{1}\colon$ $x_{1}\to x_{2}-x_{1},\quad x_{2}\to x_{2},$
$T_{1}\colon$ $x_{1}\to x_{1}+1,\quad x_{2}\to x_{2},$
$T_{2}\colon$ $x_{1}\to x_{1},\quad x_{2}\to x_{2}+1.$
Example 5.3 ($(R,\omega)=(C_{3},\omega_{3})$).
Let $R$ be the root system of type $C_{3}$, and $V=\mathbb{R}^{3}$. We take the simple roots
$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3},\quad\alpha_{3}=2e_{3}.$
Then we have the coroots
$\alpha^{\vee}_{1}=e_{1}-e_{2},\quad\alpha^{\vee}_{2}=e_{2}-e_{3},\quad\alpha^{\vee}_{3}=e_{3}.$
The fundamental weights are given by
$\omega_{1}=\alpha^{\vee}_{1}+\alpha^{\vee}_{2}+\alpha^{\vee}_{3},\quad\omega_{2}=\alpha^{\vee}_{1}+2\alpha^{\vee}_{2}+2\alpha^{\vee}_{3},\quad\omega_{3}=\alpha^{\vee}_{1}+2\alpha^{\vee}_{2}+3\alpha^{\vee}_{3}.$
Let us take $\omega=\omega_{3}$, and introduce coordinates $x_{1},x_{2},x_{3}$ of $V$ by
$\mathbf{x}=c\omega_{3}+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2}+x_{3}\alpha^{\vee}_{3}.$
Since $\kappa=2$, we have
$\lambda=\mathrm{e}^{-4\pi ic},$
then the $\lambda$-Fourier invariant polynomials are given by
$y_{1}=\lambda^{\frac{1}{2}}(\xi_{1}+\xi_{2}+\xi_{3}),\quad y_{2}=\lambda(\xi_{1}\xi_{2}+\xi_{1}\xi_{3}+\xi_{2}\xi_{3}),\quad y_{3}=\lambda^{\frac{3}{2}}\xi_{1}\xi_{2}\xi_{3},$
where
$\xi_{1}$ $=\lambda^{-\frac{1}{2}}\mathrm{e}^{2\pi ix_{1}}+\lambda^{\frac{1}{2}}\mathrm{e}^{-2\pi ix_{1}},\quad\xi_{2}=\lambda^{-\frac{1}{2}}\mathrm{e}^{2\pi i(x_{2}-x_{1})}+\lambda^{\frac{1}{2}}\mathrm{e}^{-2\pi i(x_{2}-x_{1})},$
$\xi_{3}$ $=\lambda^{-\frac{1}{2}}\mathrm{e}^{2\pi i(x_{3}-x_{2})}+\lambda^{\frac{1}{2}}\mathrm{e}^{-2\pi i(x_{3}-x_{2})}.$
The metric (,)(\ ,\ )( , ) has the form
$\left((\operatorname{d\!}x_{i},\operatorname{d\!}x_{j})\right)=\begin{pmatrix}2&-1&0\\ -1&2&-1\\ 0&-1&1\end{pmatrix}^{-1}=\begin{pmatrix}1&1&1\\ 1&2&2\\ 1&2&3\end{pmatrix}.$
In the coordinates $y_{1},y_{2},y_{3}$ we have
$\left(g_{\lambda}^{ij}\right)$ $=\frac{1}{4\pi^{2}}\left((\operatorname{d\!}y_{i},\operatorname{d\!}y_{j})\right)=\frac{1}{4\pi^{2}}\sum_{k,r=1}^{3}\frac{\partial y_{i}}{\partial x_{k}}\frac{\partial y_{j}}{\partial x_{r}}(\operatorname{d\!}x_{k},\operatorname{d\!}x_{r})$
$=\begin{pmatrix}-y_{1}^{2}+2y_{2}+12\lambda&-y_{1}y_{2}+3y_{3}+8\lambda y_{1}&-y_{1}y_{3}+4\lambda y_{2}\\ -y_{1}y_{2}+3y_{3}+8\lambda y_{1}&-2y_{2}^{2}+2y_{1}y_{3}-8\lambda y_{2}+8\lambda y_{1}^{2}&-2y_{2}y_{3}+4\lambda y_{1}y_{2}-12\lambda y_{3}\\ -y_{1}y_{3}+4\lambda y_{2}&-2y_{2}y_{3}+4\lambda y_{1}y_{2}-12\lambda y_{3}&-3y_{3}^{2}-8\lambda y_{1}y_{3}+4\lambda y_{2}^{2}\end{pmatrix}.$
It follows that $y_{1},y_{2},y_{3}$ are pencil generators, and the corresponding flat pencil is given by
$\left(\eta^{ij}\right)$ $=\begin{pmatrix}12&8y_{1}&4y_{2}\\ 8y_{1}&-8y_{2}+8y_{1}^{2}&4y_{1}y_{2}-12y_{3}\\ 4y_{2}&4y_{1}y_{2}-12y_{3}&-8y_{1}y_{3}+4y_{2}^{2}\end{pmatrix},$
$\left(g^{ij}\right)$ $=\begin{pmatrix}-y_{1}^{2}+2y_{2}&-y_{1}y_{2}+3y_{3}&-y_{1}y_{3}\\ -y_{1}y_{2}+3y_{3}&-2y_{2}^{2}+2y_{1}y_{3}&-2y_{2}y_{3}\\ -y_{1}y_{3}&-2y_{2}y_{3}&-3y_{3}^{2}\end{pmatrix}.$
The flat coordinates $t_{1},t_{2},t_{3}$ of $\eta$ are given by the polynomials
$y_{1}=2\sqrt{3}\,t_{2},\quad y_{2}=4t_{2}^{2}-4t_{1}t_{3},\quad y_{3}=\frac{8}{3\sqrt{3}}\left(t_{2}^{3}+t_{1}^{3}+t_{3}^{3}-3t_{1}t_{2}t_{3}\right).$
In these flat coordinates the metric $\eta$ has the form
$\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}.$
Thus we obtain the a generalized Frobenius manifold $\mathcal{M}(C_{3},\omega_{3})$ with potential
$F=-\frac{1}{9}t_{1}^{3}t_{2}-\frac{1}{36}t_{2}^{4}-\frac{1}{3}t_{1}t_{2}^{2}t_{3}-\frac{1}{6}t_{1}^{2}t_{3}^{2}-\frac{1}{9}t_{2}t_{3}^{3}.$
The Euler vector field and the unit vector field are given by
$E=$ $\,\frac{1}{2}t_{1}\frac{\partial}{\partial t_{1}}+\frac{1}{2}t_{2}\frac{\partial}{\partial t_{2}}+\frac{1}{2}t_{3}\frac{\partial}{\partial t_{3}},$
$e=$ $\,\frac{3}{2\left(t_{1}^{3}+t_{2}^{3}+t_{3}^{3}-3t_{1}t_{2}t_{3}\right)}$
×$\,\times\left((t_{1}t_{2}-t_{3}^{2})\frac{\partial}{\partial t_{1}}+(t_{1}t_{3}-t_{2}^{2})\frac{\partial}{\partial t_{2}}+(t_{2}t_{3}-t_{1}^{2})\frac{\partial}{\partial t_{3}}\right).$
We also have $e^{\flat}=-\frac{1}{2}\mathrm{d}\log y_{3}$.
Note that, if we take the flat coordinates $v_{1},v_{2},v_{3}$ of $\eta$ via a linear transformation
$t_{1}$ $=\frac{1}{\sqrt{3}}(v_{1}+\zeta v_{2}+\zeta^{2}v_{3}),$
$t_{2}$ $=\frac{1}{\sqrt{3}}(v_{1}+v_{2}+v_{3}),$
$t_{3}$ $=\frac{1}{\sqrt{3}}(v_{1}+\zeta^{2}v_{2}+\zeta v_{3}),$
where $\zeta=\frac{-1+\sqrt{3}i}{2}$, then in these coordinates the metric $\eta$ has the form
$\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix},$
and the metric $g$ has the form
$\left(g^{\alpha\beta}\right)=\begin{pmatrix}-v_{1}^{2}&0&0\\ 0&-v_{2}^{2}&0\\ 0&0&-v_{3}^{2}\end{pmatrix}.$
The potential of the generalized Frobenius manifold is given by
$F=-\frac{1}{12}v_{1}^{4}-\frac{1}{12}v_{2}^{4}-\frac{1}{12}v_{3}^{4},$
the Euler vector field and the unit vector field have the form
$E=\frac{1}{2}v^{1}\frac{\partial}{\partial v^{1}}+\frac{1}{2}v^{2}\frac{\partial}{\partial v^{2}}+\frac{1}{2}v^{3}\frac{\partial}{\partial v^{3}},\quad e=-\frac{1}{2v^{1}}\frac{\partial}{\partial v^{1}}-\frac{1}{2v^{2}}\frac{\partial}{\partial v^{2}}-\frac{1}{2v^{3}}\frac{\partial}{\partial v^{3}}.$
It is clear that, up to a factor, we have the isomorphism
$\mathcal{M}(C_{3},\omega_{3})~{}\cong~{}\mathcal{M}(A_{1},\omega_{1})^{\oplus 3}.$
Moreover, since the flat coordinates of $g$ has the form
$x_{1}\big{|}_{\lambda=0}=$ $\frac{1}{2\pi i}\log(2v_{1}),\quad x_{2}\big{|}_{\lambda=0}=\frac{1}{2\pi i}\left(\log(2v_{2})+\log(2v_{1})\right),$
$x_{3}\big{|}_{\lambda=0}=$ $\frac{1}{2\pi i}\left(\frac{1}{2}\log(2v_{3})+\frac{1}{2}\log(2v_{2})+\frac{1}{2}\log(2v_{1})\right),$
we know that the monodromy group is composed of only those translations on each $x_{i}$ by integer numbers. Thus we have
$\mathrm{Mono}(\mathcal{M}(C_{3},\omega_{3}))~{}\cong~{}\{e\}\rtimes\mathbb{Z}^{3}~{}<~{}\mathrm{Stab}(\omega_{3})\rtimes\mathbb{Z}^{3}.$ (5.3)
Here $\mathrm{Stab}(\omega_{3})$ is nontrivial since $\sigma_{1},\sigma_{2}\in\mathrm{Stab}(\omega_{3})$. This example shows that $\mathrm{Mono}(\mathcal{M}(R,\omega))$ is not always isomorphic to $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$.
Example 5.4 ($(R,\omega)=(B_{3},\omega_{1})$).
Let $R$ be the root system of type $B_{3}$, and $V=\mathbb{R}^{3}$. We take the simple roots
$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3},\quad\alpha_{3}=e_{3}.$
Then we have the coroots
$\alpha^{\vee}_{1}=e_{1}-e_{2},\quad\alpha^{\vee}_{2}=e_{2}-e_{3},\quad\alpha^{\vee}_{3}=2e_{3}.$
The fundamental weights are given by
$\omega_{1}=\alpha^{\vee}_{1}+\alpha^{\vee}_{2}+\frac{1}{2}\alpha^{\vee}_{3},\quad\omega_{2}=\alpha^{\vee}_{1}+2\alpha^{\vee}_{2}+\alpha^{\vee}_{3},\quad\omega_{3}=\frac{1}{2}\alpha^{\vee}_{1}+\alpha^{\vee}_{2}+\frac{3}{4}\alpha^{\vee}_{3}.$
Let us take $\omega=\omega_{1}$, and introduce coordinates $x_{1},x_{2},x_{3}$ of $V$ by
$\mathbf{x}=c\omega_{1}+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2}+x_{3}\alpha^{\vee}_{3}.$
Since $\kappa=1$, we have
$\lambda=\mathrm{e}^{-2\pi ic}.$
The $\lambda$-Fourier invariant polynomials are given by
$y_{1}=\lambda(\xi_{1}+\xi_{2}+\xi_{3}),\quad y_{2}=\lambda(\xi_{1}\xi_{2}+\xi_{1}\xi_{3}+\xi_{2}\xi_{3}),\quad y_{3}=\lambda^{\frac{1}{2}}\tilde{\xi}_{1}\tilde{\xi}_{2}\tilde{\xi}_{3}.$
Here
$\xi_{1}$ $=\lambda^{-1}\mathrm{e}^{2\pi ix_{1}}+\lambda\mathrm{e}^{-2\pi ix_{1}},\quad\xi_{2}=\mathrm{e}^{2\pi i(x_{2}-x_{1})}+\mathrm{e}^{-2\pi i(x_{2}-x_{1})},$
$\xi_{3}$ $=\mathrm{e}^{2\pi i(2x_{3}-x_{2})}+\mathrm{e}^{-2\pi i(2x_{3}-x_{2})},$
and
$\tilde{\xi}_{1}$ $=\lambda^{-\frac{1}{2}}e^{\pi ix_{1}}+\lambda^{\frac{1}{2}}\mathrm{e}^{-\pi ix_{1}},\quad\tilde{\xi}_{2}=e^{\pi i(x_{2}-x_{1})}+e^{-\pi i(x_{2}-x_{1})},$
$\tilde{\xi}_{3}$ $=e^{\pi i(2x_{3}-x_{2})}+e^{-\pi i(2x_{3}-x_{2})}.$
The metric (,)(\ ,\ )( , ) has the form
$\left((dx_{i},dx_{j})\right)=\begin{pmatrix}2&-1&0\\ -1&2&-2\\ 0&-2&4\end{pmatrix}^{-1}=\begin{pmatrix}1&1&\frac{1}{2}\\ 1&2&1\\ \frac{1}{2}&1&\frac{3}{4}\end{pmatrix}.$
In this case, $y_{1},y_{2},y_{3}$ are not pencil generators. Let us consider the proper generators with the form
$z_{1}=y_{1}+a_{1}\lambda,\quad z_{2}=y_{2}+a_{2}\lambda,\quad z_{3}=z_{3},$
here $a_{1},a_{2}$ are two constants. Then there are three possible choices of $a_{1}$ and $a_{2}$ such that $z_{1},z_{2}$ and $z_{3}$ form a set of pencil generators, they are given by
$a_{1}=6,\quad a_{2}=-12.$ (5.4)
$a_{1}=2,\quad a_{2}=4,$ (5.5)
$a_{1}=-2,\quad a_{2}=4.$ (5.6)
We first consider the $a_{1}=6,a_{2}=-12$ case. In this case we have
$\left(g_{\lambda}^{ij}\right)=$ $\,\frac{1}{4\pi^{2}}\left((\operatorname{d\!}z_{i},\operatorname{d\!}z_{j})\right)=\frac{1}{4\pi^{2}}\sum_{k,r=1}^{3}\frac{\partial z_{i}}{\partial x_{k}}\frac{\partial z_{j}}{\partial x_{r}}(\operatorname{d\!}x_{k},\operatorname{d\!}x_{r})$
$=$ $\,\begin{pmatrix}-z_{1}^{2}&-z_{1}z_{2}&-\frac{1}{2}z_{1}z_{3}\\ -z_{1}z_{2}&-2z_{2}^{2}-4z_{1}z_{2}+2z_{1}z_{3}^{2}&2z_{1}z_{3}-z_{2}z_{3}\\ -\frac{1}{2}z_{1}z_{3}&2z_{1}z_{3}-z_{2}z_{3}&-\frac{3}{4}z_{3}^{2}+4z_{1}+z_{2}\end{pmatrix}$
$\,+\lambda\begin{pmatrix}12z_{1}+2z_{2}&3z_{3}^{2}-16z_{1}&6z_{3}\\ 3z_{3}^{2}-16z_{1}&-32z_{2}-12z_{3}^{2}-64z_{1}&-24z_{3}\\ 6z_{3}&-24z_{3}&0\end{pmatrix}$
$=$ $\,\left(g^{ij}\right)+\lambda\left(\eta^{ij}\right).$
The metric $\eta$ has the flat coordinates
$t_{1}$ $=-\frac{1}{1920z_{3}^{7/3}}\left(1280z_{1}^{2}+640z_{1}z_{2}+80z_{2}^{2}-480z_{1}z_{3}^{2}+120z_{2}z_{3}^{2}-27z_{3}^{4}\right),$
$t_{2}$ $=\frac{1}{8z_{3}}(16z_{1}+4z_{2}-z_{3}^{2}),$
$t_{3}$ $=z_{3}^{1/3}.$
In these coordinates we have
$(\eta^{ij})=\begin{pmatrix}0&0&1\\ 0&3&0\\ 1&0&0\end{pmatrix}.$
We then obtain a generalized Frobenius manifold structure $\mathcal{M}(B_{3},\omega_{1},1)$ with potential
$F=$ $\,\frac{t_{1}^{2}t_{2}}{3t_{3}}-\frac{t_{1}^{2}\ t_{3}^{2}}{12}-\frac{2t_{1}t_{2}^{3}}{27t_{3}^{2}}-\frac{1}{36}\ t_{1}t_{2}^{2}t_{3}+\frac{1}{360}t_{1}t_{2}\ t_{3}^{4}-\frac{t_{1}t_{3}^{7}}{30240}$
$\,+\frac{t_{2}^{5}}{135\ t_{3}^{3}}-\frac{t_{2}^{4}}{432}+\frac{t_{2}^{3}\ t_{3}^{3}}{2160}-\frac{t_{2}^{2}t_{3}^{6}}{8640}+\frac{t_{2}\ t_{3}^{9}}{172800}-\frac{t_{3}^{12}}{7603200}.$
The Euler vector field and the unit vector field are given by
$E$ $=\frac{5}{6}t_{1}\frac{\partial}{\partial t_{1}}+\frac{1}{2}t_{2}\frac{\partial}{\partial t_{2}}+\frac{1}{6}t_{3}\frac{\partial}{\partial t_{3}},$
$e$ $=-\frac{6}{D}\left(\left(40t_{1}+20t_{2}t_{3}^{2}+t_{3}^{5}\right)\frac{\partial}{\partial t_{1}}+\left(40t_{2}+20t_{3}^{3}\right)\frac{\partial}{\partial t_{2}}+40t_{3}\frac{\partial}{\partial t_{3}}\right).$
Here $D=40t_{2}^{2}+240t_{1}t_{3}+40t_{2}t_{3}^{3}+t_{3}^{6}$. We also have $e^{\flat}=-\mathrm{d}\log z_{1}$.
Now we consider the $a_{1}=2,a_{2}=4$ case. We have
$\left(g_{\lambda}^{ij}\right)=$ $\,\frac{1}{4\pi^{2}}\left((\operatorname{d\!}z_{i},\operatorname{d\!}z_{j})\right)=\frac{1}{4\pi^{2}}\sum_{k,r=1}^{3}\frac{\partial z_{i}}{\partial x_{k}}\frac{\partial z_{j}}{\partial x_{r}}(\operatorname{d\!}x_{k},\operatorname{d\!}x_{r})$
$=$ $\,\begin{pmatrix}-z_{1}^{2}&-z_{1}z_{2}&-\frac{1}{2}z_{1}z_{3}\\ -z_{1}z_{2}&-2z_{2}^{2}-4z_{1}z_{2}+2z_{1}z_{3}^{2}&2z_{1}z_{3}-z_{2}z_{3}\\ -\frac{1}{2}z_{1}z_{3}&2z_{1}z_{3}-z_{2}z_{3}&-\frac{3}{4}z_{3}^{2}+4z_{1}+z_{2}\end{pmatrix}$
$\,+\lambda\begin{pmatrix}4z_{1}+2z_{2}&3z_{3}^{2}-4z_{2}&4z_{3}\\ 3z_{3}^{2}-4z_{2}&16z_{2}-4z_{3}^{2}&0\\ 4z_{3}&0&0\end{pmatrix}$
$=$ $\,\left(g^{ij}\right)+\lambda\left(\eta^{ij}\right).$
The metric $\eta$ has the flat coordinates
$t_{1}=\frac{48z_{1}+12z_{2}-5z_{3}^{2}}{48\sqrt{z_{3}}},\quad t_{2}=\frac{1}{4}\sqrt{4z_{2}-z_{3}^{2}},\quad t_{3}=\frac{1}{2}\sqrt{z_{3}}.$
In these coordinates we have
$\left(\eta^{ij}\right)=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}.$
We then obtain a generalized Frobenius manifold structure $\mathcal{M}(B_{3},\omega_{1},2)$ with potential
$F=$ $\,-\frac{t_{2}^{4}}{48}+\frac{t_{1}^{3}}{24t_{3}}-\frac{1}{4}t_{1}t_{2}^{2}t_{3}-\frac{t_{1}^{2}t_{3}^{2}}{12}+\frac{t_{2}^{2}t_{3}^{4}}{24}+\frac{t_{1}t_{3}^{5}}{180}-\frac{t_{3}^{8}}{2268}.$
The Euler vector field and the unit vector field are given by
$E$ $=\frac{3}{4}t_{1}\frac{\partial}{\partial t_{1}}+\frac{1}{2}t_{2}\frac{\partial}{\partial t_{2}}+\frac{1}{4}t_{3}\frac{\partial}{\partial t_{3}},$
$e$ $=\frac{1}{3t_{2}^{2}-6t_{1}t_{3}-2t_{3}^{4}}\left(\left(6t_{1}+8t_{3}^{3}\right)\frac{\partial}{\partial t_{1}}-6t_{2}\frac{\partial}{\partial t_{2}}+6t_{3}\frac{\partial}{\partial t_{3}}\right).$
We also have $e^{\flat}=-\mathrm{d}\log z_{1}$.
Finally, let us consider the $a_{1}=-2,a_{2}=4$ case. We have
$\left(g_{\lambda}^{ij}\right)=$ $\,\frac{1}{4\pi^{2}}\left((\operatorname{d\!}z_{i},\operatorname{d\!}z_{j})\right)=\frac{1}{4\pi^{2}}\sum_{k,r=1}^{3}\frac{\partial z_{i}}{\partial x_{k}}\frac{\partial z_{j}}{\partial x_{r}}(\operatorname{d\!}x_{k},\operatorname{d\!}x_{r})$
$=$ $\,\begin{pmatrix}-z_{1}^{2}&-z_{1}z_{2}&-\frac{1}{2}z_{1}z_{3}\\ -z_{1}z_{2}&-2z_{2}^{2}-4z_{1}z_{2}+2z_{1}z_{3}^{2}&2z_{1}z_{3}-z_{2}z_{3}\\ -\frac{1}{2}z_{1}z_{3}&2z_{1}z_{3}-z_{2}z_{3}&-\frac{3}{4}z_{3}^{2}+4z_{1}+z_{2}\end{pmatrix}$
$\,+\lambda\begin{pmatrix}-4z_{1}+2z_{2}&3z_{3}^{2}-8z_{2}&2z_{3}\\ 3z_{3}^{2}-8z_{2}&4z_{3}^{2}&8z_{3}\\ 2z_{3}&8z_{3}&16\end{pmatrix}$
$=$ $\,\left(g^{ij}\right)+\lambda\left(\eta^{ij}\right).$
The metric $\eta$ has the flat coordinates
$t_{1}=-\frac{48z_{1}+8z_{2}-5z_{3}^{2}}{48\left(4z_{2}-z_{3}^{2}\right)^{1/4}},\quad t_{2}=\frac{1}{4}z_{3},\quad t_{3}=\frac{1}{2}\left(4z_{2}-z_{3}^{2}\right)^{1/4}.$
In these coordinates we have
$\left(\eta^{ij}\right)=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix},$
and we obtain the same potential as the one given in the previous case with $a_{1}=2,a_{2}=4$. In all cases, the monodromy groups are given by
$\mathrm{Mono}(\mathcal{M}(B_{3},\omega_{1},k))=\mathrm{Stab}(\omega_{1})\rtimes\mathbb{Z}^{3}=\langle\sigma_{2},\sigma_{3}\rangle\rtimes\mathbb{Z}^{3}\cong D_{4}\rtimes\mathbb{Z}^{3},\quad k=1,2,3.$
Example 5.5 ($(R,\omega)=(G_{2},\omega_{2})$).
Let $R$ be the root system of type $G_{2}$, and $V$ be the subspace of $\mathbb{R}^{3}$ spanned by the simple roots
$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=-2e_{1}+e_{2}+e_{3}.$
We have the coroots
$\alpha^{\vee}_{1}=\alpha_{1},\quad\alpha^{\vee}_{2}=\frac{1}{3}\alpha_{2}.$
The fundamental weights are given by
$\omega_{1}=2\alpha^{\vee}_{1}+3\alpha^{\vee}_{2},\quad\omega_{2}=3\alpha^{\vee}_{1}+6\alpha^{\vee}_{2}.$
Let us take $\omega=\omega_{2}$ and introduce coordinates $x_{1},x_{2}$ of $V$ by
$\mathbf{x}=c\,\omega_{2}+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2}=(x_{1}+3c)\alpha^{\vee}_{1}+(x_{2}+6c)\alpha^{\vee}_{2},$
Since $\kappa=3$, we have
$\lambda=\mathrm{e}^{-6\pi ic}.$
The $\lambda$-Fourier invariant polynomials are given by
$y_{1}=$ $\,\mathrm{e}^{-6\pi ic}\left(\mathrm{e}^{2\pi i(x_{1}+3c)}+\mathrm{e}^{-2\pi i(x_{1}+3c)}+\mathrm{e}^{2\pi i(2x_{1}-x_{2})}+\mathrm{e}^{-2\pi i(2x_{1}-x_{2})}\right.$
$\,\left.+\mathrm{e}^{2\pi i(x_{1}-x_{2}-3c)}+\mathrm{e}^{-2\pi i(x_{1}-x_{2}-3c)}\right)$
$=$ $\,\mathrm{e}^{2\pi ix_{1}}+\lambda^{2}\mathrm{e}^{-2\pi ix_{1}}+\lambda\mathrm{e}^{2\pi i(2x_{1}-x_{2})}+\lambda\mathrm{e}^{-2\pi i(2x_{1}-x_{2})}$
$+\lambda^{2}\mathrm{e}^{2\pi i(x_{1}-x_{2})}+\mathrm{e}^{-2\pi i(x_{1}-x_{2})},$
$y_{2}=$ $\,e^{-12\pi ic}\left(\mathrm{e}^{2\pi i(x_{2}+6c)}+\mathrm{e}^{-2\pi i(x_{2}+6c)}+\mathrm{e}^{2\pi i(3x_{1}-2x_{2}-3c)}+\mathrm{e}^{-2\pi i(3x_{1}-2x_{2}-3c)}\right.$
$\,\left.+\mathrm{e}^{2\pi i(x_{2}-3x_{1}-3c)}+\mathrm{e}^{-2\pi i(x_{2}-3x_{1}-3c)}\right)$
$=$ $\,\mathrm{e}^{2\pi ix_{2}}+\lambda^{4}\mathrm{e}^{-2\pi ix_{2}}+\lambda^{3}\mathrm{e}^{2\pi i(3x_{1}-2x_{2})}+\lambda\mathrm{e}^{-2\pi i(3x_{1}-2x_{2})}$
$\,+\lambda^{3}\mathrm{e}^{2\pi i(x_{2}-3x_{1})}+\lambda\mathrm{e}^{-2\pi i(x_{2}-3x_{1})}.$
The metric (,)(\ ,\ )( , ) has the form
$\left((\operatorname{d\!}x_{i},\operatorname{d\!}x_{j})\right)=\begin{pmatrix}2&-1\\ -1&\frac{2}{3}\end{pmatrix}^{-1}=\begin{pmatrix}2&3\\[3.0pt] 3&6\end{pmatrix}.$
In the $y_{1},y_{2}$ coordinates the metric does not depend at most linearly on $\lambda$, hence $y_{1}$ and $y_{2}$ are not pencil generators. Proper generators $z_{1}$ and $z_{2}$ may be introduced by
$z_{1}=y_{1}-6\lambda,\quad z_{2}=y_{2}-3\lambda y_{1}+12\lambda^{2}.$
Then in the $z_{1},z_{2}$ coordinates we have
$\left(g^{\alpha\beta}_{\lambda}\right)=\left(g^{\alpha\beta}\right)+\lambda\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}-2z_{1}^{2}+2z_{2}&-3z_{1}z_{2}\\ -3z_{1}z_{2}&-6z_{2}^{2}\end{pmatrix}+\lambda\begin{pmatrix}-12z_{1}&3z_{1}^{2}-36z_{2}\\ 3z_{1}^{2}-36z_{2}&6z_{1}^{3}-36z_{1}z_{2}\end{pmatrix}$
The metric $\eta$ has flat coordinates
$t_{1}=\frac{z_{1}}{2\left(4z_{2}-z_{1}^{2}\right)^{\frac{1}{6}}},\quad t_{2}=-\frac{1}{3}\left(4z_{2}-z_{1}^{2}\right)^{\frac{1}{6}},$
in these coordinates the metric $\eta$ has the form
$\left(\eta^{\alpha\beta}\right)=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}.$
We thus obtain a generalized Frobenius manifold with potential
$F=\frac{t_{1}^{4}}{2916t_{2}^{2}}-\frac{1}{8}t_{1}^{2}t_{2}^{2}+\frac{81t_{2}^{6}}{320}.$
The Euler vector field and the unit vector field are given by
$E$ $=\frac{2}{3}t_{1}\frac{\partial}{\partial t_{1}}+\frac{1}{3}t_{2}\frac{\partial}{\partial t_{1}},$
$e$ $=-\frac{2}{3t_{2}\left(4t_{1}^{2}+81t_{2}^{4}\right)}\left((4t_{1}^{2}+243t_{2}^{4})\frac{\partial}{\partial t_{1}}+4t_{1}t_{2}\frac{\partial}{\partial t_{2}}\right).$
In this case, $\mathrm{Stab}(\omega_{1})=\langle\sigma_{1}\rangle$, where
$\sigma_{1}(x_{1},x_{2})=(-x_{1}+x_{2},x_{2}).$
However, one has
$t_{1}=\frac{z_{1}}{2\sqrt[6]{-1}\left(\mathrm{e}^{2\pi ix_{1}}-\mathrm{e}^{-2\pi i(x_{1}-x_{2})}\right)^{\frac{1}{3}}},\quad t_{2}=-\frac{1}{3}\sqrt[6]{-1}\left(\mathrm{e}^{2\pi ix_{1}}-\mathrm{e}^{-2\pi i(x_{1}-x_{2})}\right)^{\frac{1}{3}}.$
Hence the actions of $\sigma_{1}$ on $t_{1}$ and $t_{2}$ is not trivial. In fact,
[…]
Here the value of $3]{-1}$ depends on the choice of the monodromy region of $f(z)=z^{1/3}$. Therefore we have
$\mathrm{Mono}(\mathcal{M}(G_{2},\omega_{2}))=\{e\}\rtimes\mathbb{Z}^{2}.$
Example 5.6 ($(R,\omega)=(A_{3},\omega_{1}+\omega_{3})$).
Let $R$ be the root system of type $A_{3}$, $V$ be the subspace of $\mathbb{R}^{4}$ spanned by the simple roots
$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3},\quad\alpha_{3}=e_{3}-e_{4}.$
We have $\alpha^{\vee}_{i}=\alpha_{i},i=1,2,3$, and the fundamental weights are given by
$\omega_{1}=\frac{3}{4}\alpha^{\vee}_{1}+\frac{1}{2}\alpha^{\vee}_{2}+\frac{1}{4}\alpha^{\vee}_{3},\quad\omega_{2}=\frac{1}{2}\alpha^{\vee}_{1}+\alpha^{\vee}_{2}+\frac{1}{2}\alpha^{\vee}_{3},\quad\omega_{3}=\frac{1}{4}\alpha^{\vee}_{1}+\frac{1}{2}\alpha^{\vee}_{2}+\frac{3}{4}\alpha^{\vee}_{3}.$
Let us take $\omega=\omega_{1}+\omega_{3}$, and introduce coordinates $x_{1},x_{2},x_{3}$ of $V$ by
$\mathbf{x}=c(\omega_{1}+\omega_{3})+x_{1}\alpha^{\vee}_{1}+x_{2}\alpha^{\vee}_{2}+x_{3}\alpha^{\vee}_{3},\quad\mathbf{x}\in V.$
Since $\kappa=1$, we have
$\lambda=\mathrm{e}^{-2\pi ic}.$
We have the $\lambda$-invariant Fourier polynomials
$y_{1}=$ $\,\mathrm{e}^{2\pi ix_{1}}+\lambda\mathrm{e}^{-2\pi i(x_{1}-x_{2})}+\lambda^{2}\mathrm{e}^{-2\pi ix_{3}}+\lambda\mathrm{e}^{-2\pi i(x_{2}-x_{3})},$
$y_{2}=$ $\,\mathrm{e}^{2\pi ix_{2}}+\lambda^{2}\mathrm{e}^{-2\pi ix_{2}}+\lambda\mathrm{e}^{2\pi i(x_{1}-x_{3})}+\lambda^{2}\mathrm{e}^{-2\pi i(x_{1}-x_{2}+x_{3})}$
$\,+\lambda\mathrm{e}^{-2\pi i(x_{1}-x_{3})}+\mathrm{e}^{2\pi i(x_{1}-x_{2}+x_{3})},$
$y_{3}=$ $\,\lambda^{2}\mathrm{e}^{-2\pi ix_{1}}+\lambda\mathrm{e}^{2\pi i(x_{1}-x_{2})}+\mathrm{e}^{2\pi ix_{3}}+\lambda\mathrm{e}^{2\pi i(x_{2}-x_{3})}.$
The metric on $V$ that is induced from $\mathbb{R}^{4}$ is given by
$\left((\alpha^{\vee}_{i},\alpha^{\vee}_{j})\right)=\begin{pmatrix}2&-1&0\\ -1&2&-1\\ 0&-1&2\end{pmatrix},\quad\left(\operatorname{d\!}x_{i},\operatorname{d\!}x_{j}\right)={\left((\alpha^{\vee}_{i},\alpha^{\vee}_{j})\right)}^{-1}=\frac{1}{4}\begin{pmatrix}3&2&1\\ 2&4&2\\ 1&2&3\end{pmatrix}.$
In the coordinates $y_{1},y_{2},y_{3}$ we have
$\left(g_{\lambda}^{ij}\right)$ $=\frac{1}{4\pi^{2}}\left((\operatorname{d\!}y_{i},\operatorname{d\!}y_{j})\right)=\frac{1}{4\pi^{2}}\sum_{k,r=1}^{3}\frac{\partial y_{i}}{\partial x_{k}}\frac{\partial y_{j}}{\partial x_{r}}(\operatorname{d\!}x_{k},\operatorname{d\!}x_{r})$
$=\begin{pmatrix}-\frac{3}{4}y_{1}^{2}+2\lambda y_{2}&-\frac{1}{2}y_{1}y_{2}+3\lambda y_{3}&-\frac{1}{4}y_{1}y_{3}+4\lambda^{2}\\ -\frac{1}{2}y_{1}y_{2}+3\lambda y_{3}&-y_{2}^{2}+2y_{1}y_{3}+4\lambda^{2}&-\frac{1}{2}y_{2}y_{3}+3\lambda y_{1}\\ -\frac{1}{4}y_{1}y_{3}+4\lambda^{2}&-\frac{1}{2}y_{2}y_{3}+3\lambda y_{1}&-\frac{3}{4}y_{3}^{2}+2\lambda y_{2}\end{pmatrix}.$
We introduce the pencil generators
$z_{1}=y_{1}+4\lambda,\quad z_{2}=y_{2}-6\lambda,\quad z_{3}=y_{3}+4\lambda.$
In the new coordinates, we obtain a flat pencil of metrics $\left(g^{ij}\right)+\lambda\left(\eta^{ij}\right)$ with
$\left(\eta^{ij}\right)=\begin{pmatrix}6z_{1}+2z_{2}&-3z_{1}+2z_{2}+3z_{3}&z_{1}+z_{3}\\ -3z_{1}+2z_{2}+3z_{3}&-8z_{1}-12z_{2}-8z_{3}&3z_{1}+2z_{2}-3z_{3}\\ z_{1}+z_{3}&3z_{1}+2z_{2}-3z_{3}&2z_{2}+6z_{3}\end{pmatrix}.$
The metric $\eta$ has the flat coordinates $t_{1},t_{2},t_{3}$ which are given by
$z_{1}$ $=\frac{1}{6}t_{1}^{4}+t_{1}^{2}t_{2}+\frac{1}{2}t_{2}^{2}+t_{1}t_{3},$
$z_{2}$ $=\frac{2}{3}t_{1}^{4}-t_{2}^{2}-2t_{1}t_{3},$
$z_{3}$ $=\frac{1}{6}t_{1}^{4}-t_{1}^{2}t_{2}+\frac{1}{2}t_{2}^{2}+t_{1}t_{3};$
in these coordinates the metrics $\eta$ has the form
$\left(\eta^{ij}\right)$ $=\begin{pmatrix}0&0&1\\ 0&1&0\\ 1&0&0\end{pmatrix}.$
Thus we obtain a generalized Frobenius manifold $\mathcal{M}(A_{3},\omega_{1}+\omega_{3})$ with potential
$F=$ $-\frac{t_{1}^{8}}{18144}+\frac{t_{1}^{5}\ t_{3}}{720}-\frac{t_{2}^{6}}{96t_{1}^{4}}-\frac{t_{1}^{4}\ t_{2}^{2}}{288}+\frac{t_{2}^{4}t_{3}}{16\ t_{1}^{3}}-\frac{t_{2}^{2}t_{3}^{2}}{8t_{1}^{2}}$
$-\frac{t_{1}^{2}\ t_{3}^{2}}{24}-\frac{1}{24}t_{1}t_{2}^{2}\ t_{3}+\frac{t_{3}^{3}}{24t_{1}}-\frac{t_{2}^{4}}{96},$
the Euler vector field
$E=\frac{1}{4}t_{1}\frac{\partial}{\partial t_{1}}+\frac{1}{2}t_{2}\frac{\partial}{\partial t_{2}}+\frac{3}{4}t_{3}\frac{\partial}{\partial t_{3}},$
and the unit vector field
$e=$ $\,\frac{1}{D}\left(\left(-12t_{1}^{5}-72t_{3}t_{1}^{2}-36t_{2}^{2}t_{1}\right)\frac{\partial}{\partial t_{1}}+\left(60t_{2}t_{1}^{4}-72t_{2}t_{3}t_{1}-36t_{2}^{3}\right)\frac{\partial}{\partial t_{2}}\right.$
$\,\left.+\left(-8t_{1}^{7}-60t_{3}t_{1}^{4}+120t_{2}^{2}t_{1}^{3}-72t_{3}^{2}t_{1}-36t_{2}^{2}t_{3}\right)\frac{\partial}{\partial t_{3}}\right),$
where
$D=\left(t_{1}^{4}-6t_{2}t_{1}^{2}+6t_{3}t_{1}+3t_{2}^{2}\right)\left(t_{1}^{4}+6t_{2}\ t_{1}^{2}+6t_{3}t_{1}+3t_{2}^{2}\right).$
We also have $e^{\flat}=-(\mathrm{d}\log z_{1}+\mathrm{d}\log z_{3})$. The monodromy group is given by
$\mathrm{Mono}(\mathcal{M}(A_{3},\omega_{1}+\omega_{3}))=\mathrm{Stab}(\omega_{1}+\omega_{3})=\langle\sigma_{2}\rangle\rtimes\mathbb{Z}^{3}\cong\mathbb{Z}_{2}\rtimes\mathbb{Z}^{2}.$
6. Conclusions
Given a root system $R$, a weight $\omega$ and a set of pencil generators $z^{1},\cdots,z^{\ell}$, we constructed a generalized Frobenius manifold structure $\mathcal{M}(R,\omega)$ on the orbit spaces of the associated affine Weyl groups $W_{a}(R)$. We proved that $z^{1},\cdots,z^{\ell}$ depend polynomially on the flat coordinates $t^{1},\cdots,t^{\ell}$ of the generalized Frobenius manifold, we show that the monodromy group $\mathrm{Mono}(\mathcal{M}(R,\omega))$ of the generalized Frobenius manifold can be embedded into the subgroup $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$ of $W_{a}(R)$.
In the subsequent paper [20], we will explicitly construct the pencil generators for affine Weyl groups of types $A_{\ell}$, $B_{\ell}$, $C_{\ell}$ and $D_{\ell}$ together with an appropriately chosen weight $\omega$, and compute the corresponding generalized Frobenius manifold structures in detail. For the case $(A_{\ell},\omega_{\ell})$, the generalized Frobenius manifold structure that we will present in [20] coincides with the one constructed in [7], where this generalized Frobenius manifold structure is constructed in terms of a super-potential given by a Fourier polynomial. It is shown in [7] that the bihamiltonian structure that is given by the associated flat pencil of metrics coincides with the dispersionless limit of the bihamiltonian structure of the q-deformed Gelfand-Dickey hierarchy [16, 17].
Finally, let us outline some potential directions for further research.
  1. i)
    General construction of pencil generators: Can pencil generators be constructed for every affine Weyl group $W_{a}(R)$ and an arbitrary weight $\omega$?
  2. ii)
    Precise description of monodromy groups: While we have shown that $\mathrm{Mono}(\mathcal{M}(R,\omega))$ embeds into $\mathrm{Stab}(\omega)\rtimes\mathbb{Z}^{\ell}$, can this subgroup be characterized more explicitly?
  3. iii)
    Relation to other Frobenius manifold constructions: How do these generalized Frobenius manifolds relate to earlier constructions on the orbit spaces of the finite and the extended affine Weyl groups?
  4. iv)
    Almost duality and geometric interpretations: What is the Dubrovin almost-duality structure [11] for $\mathcal{M}(R,\omega)$?
  5. v)
    Legendre transformations: What are the properties of Legendre transformations [23, 27] for these generalized Frobenius manifolds?
  6. vi)
    Associated integrable systems: Which integrable hierarchies already known in the theory of solitons are given by the topological deformations of the Principal Hiearchies [21, 22] of these generalized Frobenius manifolds?